### v1

# RefDrone: A Challenging Benchmark for Referring Expression Comprehension in Drone Scenes

Abstract
Drones have become prevalent robotic platforms with diverse applications, showing significant potential in Embodied Artificial Intelligence (Embodied AI). Referring Expression Comprehension (REC) enables drones to locate objects based on natural language expressions, a crucial capability for Embodied AI. Despite advances in REC for ground-level scenes, aerial views introduce unique challenges including varying viewpoints, occlusions and scale variations. To address this gap, we introduce RefDrone, a REC benchmark for drone scenes. RefDrone reveals three key challenges in REC: 1) multi-scale and small-scale target detection; 2) multi-target and no-target samples; 3) complex environment with rich contextual expressions. To efficiently construct this dataset, we develop RDAgent (referring drone annotation framework with multi-agent system), a semi-automated annotation tool for REC tasks. RDAgent ensures high-quality contextual expressions and reduces annotation cost. Furthermore, we propose Number GroundingDINO (NGDINO), a novel method designed to handle multi-target and no-target cases. NGDINO explicitly learns and utilizes the number of objects referred to in the expression. Comprehensive experiments with state-of-the-art REC methods demonstrate that NGDINO achieves superior performance on both the proposed RefDrone and the existing gRefCOCO datasets. The dataset and code will be publicly at https://github.com/sunzc-sunny/refdrone.
1 Introduction
Figure 1: Examples of the various challenges in RefDrone dataset.
Table 1: Comparison of referring expression comprehension datasets: RefCOCO/+/g [45, 29], gRefCOCO [22], D3 [42], RIS-CQ [13], RSVG [47], and RefDrone (ours). Avg. object: average number of objects per expression. Avg. length: words per expression. No target: expressions without referred objects. Expression type: expression generation method. Small target: small-scale target detection challenge.
RefCOCO/+/g [45, 29] gRefCOCO [22] D3 [42] RIS-CQ [13] RSVG [47] RefDrone Image source COCO [20] COCO [20] COCO [20] VG [16]+COCO [20] DIOR [18] VisDrone [49] Avg. object 1.0 1.4 1.3 3.6 2.2 3.8 Avg. length 3.6/3.5/8.4 4.9 6.3 13.2 7.5 9.0 No target Expression type Manual Manual Manual LLM (GPT3.5-t) Templated LMM (GPT4-o) Small target
Drones/UAVs have become increasingly popular in our daily lives, serving both personal and professional purposes, such as entertainment, package delivery, traffic surveillance, and emergency rescue. Their ability to move freely and their broad observation capabilities make them important platforms for Embodied AI [6, 27, 17, 8, 25]. A crucial capability in Embodied AI is Referring Expression Comprehension (REC) [35, 40, 2, 39], which serves as a critical bridge between natural language understanding and visual perception. REC requires drones to localize specific objects in images based on natural language expressions. However, existing REC datasets primarily focus on ground-level perspectives, such as the RefCOCO [45, 29] dataset. The application of REC in drone-based scenarios presents unique challenges, including extreme viewpoint variations, occlusions, and scale variations across objects.
In this work, we introduce RefDrone, a challenging benchmark designed for REC in drone scenes. The RefDrone dataset contains 17,900 referring expressions across 8,536 images, comprising 63,679 object instances. As illustrated in Figure 1, RefDrone poses three key challenges that distinguish it from existing datasets: (1) multi-scale and small-scale target detection, with 31% small objects and 14% large objects; (2) multi-target and no-target samples, where expressions can refer to any number of objects from 0 to 242; and (3) complex environment with rich contextual expressions. The environmental complexity manifests in varied viewpoints, diverse lighting conditions, and complex backgrounds. The contextual richness is reflected in expressions that describe spatial relations, attributes, and interactive relationships.
To efficiently construct this comprehensive dataset, we develop RDAgent (referring drone annotation framework with multi-agent system), a semi-automated annotation pipeline. RDAgent restructures traditional annotation workflows into an interactive system of multiple agents and human annotators. In this framework, agents with diverse roles collaborate through feedback loops to generate and verify annotations, while human involvement is minimized to quality control and minor adjustments. RDAgent significantly reduces annotation costs while maintaining high-quality standards for complex expressions, and it can be extended to generate annotations for other REC tasks.
Furthermore, we propose a novel method called Number GroundingDINO (NGDINO). Our key insight is that utilizing the number information of referred objects enhances the handling of multi-target and no-target samples. NGDINO comprises three key components: (1) a number prediction head that estimates target object quantities, (2) a set of learnable number-queries encoding numerical patterns across different quantity samples, and (3) a number cross-attention module that integrates number queries with detection queries. Extensive experiments on both our drone-specific RefDrone benchmark and the general-domain gRefCOCO [22] dataset reveal that leveraging the number information greatly improves performance on multi-target and no-target samples.
In summary, our contributions are listed as follows:
  1. 1.
    RefDrone Benchmark: The first comprehensive benchmark for referring expression comprehension in drone scenes. RefDrone poses three key challenges and provides comprehensive baseline evaluations across both specialized models and large multimodal models.
  2. 2.
    RDAgent Annotation Framework: A novel semi-automated annotation framework employs a multi-agent system. RDAgent reduces annotation costs while ensuring high-quality referring expressions.
  3. 3.
    NGDINO Method: A novel method specifically designed to handle multi-target and no-target samples through three components: a number prediction head, learnable number-queries, and number cross-attention. NGDINO achieves superior performance on both RefDrone and gRefCOCO datasets.
2 Related works
2.1 Referring expression understanding datasets
Referring expression understanding identifies specific regions within images or videos using natural language expressions. Two primary subtasks are Referring Expression Comprehension (REC) and Referring Expression Segmentation (RES). REC outputs detection bounding boxes, while RES outputs segmentation masks. Early datasets such as ReferIt [15] and RefCOCO [45] are pioneering but are limited to single-target expressions and simple expressions. Subsequently, datasets [22, 37, 13, 42] introduce more challenges. gRefCOCO [22] supports multi-target expressions. OV-VG [37] enables an open-vocabulary setting. D3 [42] and RIS-CQ [13] introduce more complex expressions. Additionally, domain-specific datasets [47, 30, 32] emerge to address unique applications. RSVG [47] focuses on remote sensing, RAVAR [30] targets video action recognition, and RIO [32] addresses affordance detection. However, drone scenes remain unexplored in REC tasks, which motivates our introduction of the RefDrone benchmark.
Recent advances in dataset creation have been significantly influenced by large language models (LLMs). Notable works such as LLaVA [24] and Ferret [44] show the effectiveness of LLMs in generating instruction tuning data. Similarly, RIS-CQ [13] and RIO [32] employ LLMs to generate complex referring expressions. However, these approaches have two limitations. First, LLMs are used mainly as text generators, potentially producing expressions without proper grounding in visual context. Second, the lack of iterative feedback mechanisms in the annotation pipelines may degrade annotation quality.
Figure 2: The overview of the RefDrone annotation process with RDAgent. Multiple specialized agents collaborate both with each other and human annotators through iterative feedback loops to generate high-quality annotations.
2.2 Referring expression comprehension methods
Referring expression comprehension methods can be broadly categorized into specialist models and large multimodal models (LMMs). Specialist models include two-stage and one-stage methods. Two-stage methods [12, 23, 10, 36, 9] typically approach REC as a ranking task, first generating region proposals and then ranking them based on language input. However, they are often criticized for slow inference speeds. In contrast, one-stage methods [14, 19, 26, 7, 43, 28] directly predict target regions guided by language. These approaches leverage transformer architectures to enable cross-modal interactions between visual and textual features. Among these methods, GroundingDINO (GDINO) [28] has gained widespread attention for its impressive results in both open-set object detection and REC tasks. However, GDINO lacks explicit mechanisms to handle multi-target or no-target scenarios.
Recent studies have increasingly evaluated LMMs [24, 44, 3, 5, 1, 31, 38, 21, 48, 41] on REC tasks to assess their visual-language understanding capabilities. These models leverage large-scale referring instruction tuning data to achieve remarkable performance. Despite their broad capabilities, LMMs consistently struggle with small object detection, primarily due to their inherent input resolution constraints. This constraint forces image downsampling, resulting in the loss of fine-grained visual details.
3 RefDrone benchmark
3.1 Data source
The RefDrone dataset builds upon VisDrone2019-DET [49], a drone-captured dataset for object detection. The source images are collected across multiple scenarios, illumination conditions, and flying altitudes. To ensure meaningful visual content, we implement two filtering criteria. At the image level, we exclude images containing less than 3 objects. At the object level, we exclude objects with bounding box areas below 64 pixels. The original VisDrone2019-DET annotations provide object categories and bounding box coordinates. We convert the bounding box to a normalized center point (range 0-1). This approach reduces the input context for LMMs while preserving spatial information.
3.2 RDAgent for semi-automated annotation
RDAgent (referring drone annotation framework with multi-agent system) is a semi-automated annotation framework that enhances traditional annotation pipelines by integrating LMMs with human annotators. Central to RDAgent are multiple LMM agents, specifically GPT-4o, each perform specialized tasks via distinct system prompts. As shown in Figure 2, RDAgent employs five structured steps.
Step 1: scene understanding. A GPT-4o-based captioning agent generates three diverse textual descriptions per image. These captions provide contextual perspectives (e.g., object relationships, spatial layouts) to anchor subsequent referring expression generation.
Figure 3: Object size distribution analysis. (a) Object size distribution in RefDrone dataset (small: $<32^{2}=1024$ pixels, normal: 1024 to 9216 pixels, large: $>96^{2}=9216$ pixels). (b) Object size histograms in RefDrone and gRefCOCO [22] datasets.
Step 2: color categorization. Color attributes serve as crucial discriminative features in referring expressions. Color attributes are extracted using a hybrid pipeline: (1) a WideResNet-101 [46] model trained on HSV color space-based labels, refined via manual validation, and (2) a dedicated LMM agent that verifies predictions to mitigate illumination- or occlusion-induced errors.
Step 3: expression generation. To address the complexity in referring expression generation, we reformulate the referring expression generation task as an object grouping problem. The target of the agent is to group semantically related objects and to provide appropriate reasons for each group. The reasons serve as the referring expressions. Furthermore, a dynamic feedback loop triggers color categorization (Step 2) when novel colors are detected, ensuring color attribute consistency.
Step 4: quality evaluation. An evaluator agent assesses the annotation semantic accuracy and referential uniqueness of object-expression pairs, categorizing annotations as:
  • “Yes.” The annotation is accurate and unique to the described objects.
  • “No.” The annotation is inaccurate or non-unique, with a detailed explanation of the discrepancy.
Annotations marked “Yes” proceed to Step 5. Annotations marked “No” activate specific feedback: expressions with semantic issues are returned to Step 3 (expression generation), while color-related inaccuracies are sent to Step 2 (color categorization) for refinement.
Step 5: human verification. Human annotators review annotation outputs in three tiers:
  • Direct acceptance. Annotations satisfying all quality criteria are approved for the final dataset.
  • Refinement required. Annotations with minor errors are corrected through human editing.
  • Significant issues. Annotations with major errors or unclear descriptions are marked for complete reworking.
Annotations with significant issues enter a feedback loop, returning to Step 4 (quality evaluation). If an annotation repeatedly fails to meet standards, it becomes a no-target sample. These feedback refinement loops ensure that all annotations meet our quality standards and that no-target expressions still relate to the image content.
Each step utilizes LMMs through in-context learning with manually designed prompts and examples (see Supplementary Material). RDAgent requires only \$0.0539 pre expression in GPT4-o API cost and reduces human annotation effort by 88% compared to fully manual annotations. This cost-performance balance makes the framework scalable for large-scale datasets.
Figure 4: Object number distribution per expression in gRefCOCO [22] and RefDrone datasets.
Figure 5: Word frequency visualization in RefDrone dataset.
3.3 Dataset analysis
The RefDrone dataset contains 17,900 referring expressions across 8,536 images, comprising 63,679 object instances in 10 categories. We maintain the original train, validation, and test splits from VisDrone2019-DET [49]. The dataset features an average expression length of 9.0 words, with each expression referring to an average of 3.8 objects. Figure 1 gives several examples that relate to the three key features and challenges in RefDrone:
1) Multi-scale and small-scale target detection. Figure 3 presents an analysis of object scale distribution in RefDrone, with (a) showing the overall distribution and (b) offering a comparison with the gRefCOCO [22] dataset. In the RefDrone dataset, small-scale objects constitute 31% of all instances, while large-scale objects account for 14%. The detailed scale distribution in RefDrone demonstrates greater variance compared to gRefCOCO, highlighting the multi-scale challenges in RefDrone.
2) Multi-target and no-target samples. The RefDrone dataset includes 11,362 multi-target and 847 no-target expressions, with the number of referred targets ranging from 0 to 242. Figure 4 illustrates the target number distribution, revealing higher complexity in multi-target scenarios compared to gRefCOCO, where expressions typically refer to no more than two objects.
3) Complex environment with rich contextual expressions. The images present inherent complexity including diverse viewpoints, varying lighting conditions, and complex background environments. The referring expressions go beyond simple object attributes (e.g., color, category) and spatial relationships (e.g., ‘left’, ‘near’), incorporating rich object-object interactions (e.g., ‘the white trucks carrying livestock’) and object-environment interactions (e.g., ‘the white cars line up at the intersection’). Figure 5 presents word cloud visualizations of (a) complete expressions and (b) background terms, highlighting the diverse vocabulary used in descriptions.
3.4 Dataset comparison
Table 1 compares RefDrone with existing REC datasets. RefDrone stands out by its average number of referred objects and its use of LMM for expression generation. The expressions offer richer contextual details compared to template-based or human-annotated expressions. While RIS-CQ [13] employs LLM for expression generation, it lacks visual content during the generation process. This often results in expressions that, despite being linguistically complex, may be disconnected from visual content. Similarly, RSVG [47], although it focuses on small targets, falls short in terms of expression quality. In contrast, RefDrone comprehensively presents three challenges, establishing it as a challenging benchmark in REC tasks.
3.5 Evaluation metrics
We introduce instance-level metrics extending traditional REC metrics to address multi-target challenges. Previous benchmarks primarily focus on image-level metrics, which are sufficient for single-target or few-target samples. However, these metrics fail to handle expressions referring to multiple target objects—potentially scaling up to 100 instances per expression in our task. To address this limitation, we introduce instance-level metrics that provide a more granular evaluation of a model’s capability to identify individual target objects.
Instance-level metrics: Accinst. and F1inst. measure the accuracy of individual bounding box predictions. We compute the intersection over union (IoU) between the true bounding boxes and the predicted bounding boxes. An IoU $\geq$ 0.5 is considered a true positive (TP), otherwise it is a false positive (FP). Unmatched true bounding boxes are counted as false negatives (FN). For no-target samples, a prediction without any bounding box is a true negative (TN), otherwise it is a false positive (FP). We calculate:
$\text{Acc}_{inst.}=\frac{TP+TN}{TP+TN+FP+FN}$ (1)
and
[…]
Image-level metrics: Accimg. and F1img. evaluate the overall accuracy of predictions per image. An image is considered a true positive (TP) if all predictions match the true bounding boxes, otherwise it is a false positive (FP). For no-target samples, a prediction without any bounding box is a true negative (TN), otherwise it is a false positive (FP). We calculate Accimg. and F1img. using the same formulas as Accinst. and F1inst., respectively.
4 Proposed method
We introduce Number GroundingDINO (NGDINO) to address multi-target and no-target challenges based on GDINO [28]. Our insight is that utilizing number information of referred objects enhances model performance on these challenges. As illustrated in Figure 6, we introduce three components: (1) a number prediction head, (2) number-queries, and (3) a number cross-attention module.
Figure 6: Architecture of a single decoder layer in Number GroundingDINO. Key modifications from GDINO [28] (highlighted in yellow box) include: (1) number prediction head (FFN) to predict target number, (2) number-queries selected through the predicted number, (3) number cross-attention between selected number-queries and detection-queries.
4.1 NGDINO
Following GDINO, our model employs a dual-encoder-single-decoder architecture. The framework consists of an image backbone for visual feature extraction, a text backbone for text feature extraction, a feature enhancer for cross-modal fusion, a language-guided query selection module for detection-query initialization, and a cross-modality decoder for box refinement. Within decoder layers, detection-queries are fed into a self-attention layer, an image cross-attention layer to combine image features, a text cross-attention layer to combine text features, and an FFN layer. We mainly improve the decoder part, highlighted by the yellow box in Figure 6.
The number prediction head, structured similarly to the detection head with an FFN layer, predicts the number of referred objects from detection-queries. These detection-queries are used to predict the objects, containing number information. The number-queries are learnable embeddings designed to capture various numerical patterns. These queries are initialized randomly with dimensions (bs, length_nquery, ndim), where bs is the batch size, length_nquery represents the number of query lengths, and ndim denotes the feature dimension. The predicted number guides the selection of number-queries through the query selection in Algorithm 1. The selected number-queries have a fixed length of length_snquery.
We implement number cross-attention in parallel with self-attention to integrate the number information. In the number cross-attention, the selected number-queries serve as keys and values, while detection-queries serve as queries. The outputs are added to the self-attention features.
4.2 Loss function
[…]
""" Input: detection_queries:(bs,length_dquery,ndim) number_queries:(bs,length_nquery,ndim) Output: selected_number_queries:(bs,length_snquery,ndim) """ number_feat = FFN(detection_queries) # (bs, length_dquery, 1) number_feat = number_feat.mean(dim=1) # (bs, 1) number = sigmoid(number_feat) # (bs, 1) index = slice(length_snquery*number, length_snquery*(number+1)) selected_number_queries = number_queries[index] # (bs, length_snquery, ndim)
Algorithm 1 Query selection.
5 Experiments
5.1 Construction of RDAgent and NGDINO
RDAgent. A straightforward method is to adapt RDAgent for REC tasks. This adaptation automates the process by replacing the object input with predictions from an object detector and removing the human verification step. For a fair comparison, we employ Faster R-CNN [34] as the object detector, which is trained on the VisDrone2019-DET [49] dataset. The agent component utilizes GPT-4o [^1] https://openai.com/index/hello-gpt-4o/.
NGDINO. The NGDINO adopts a two-stage procedure to maintain training stability. First, we pre-train the number prediction head on the RefDrone dataset while initializing other components with GDINO [28] parameters. Then, we fine-tune the entire model.
5.2 Implementation details
We establish the benchmark with 13 representative models that can perform REC tasks, comprising 3 specialist models and 10 LMMs. The specialist models include MDETR [14], GLIP [19], and GDINO [28]. For LMMs, we evaluate Shikra [5], ONE-PEACE [38], SPHINX-v2 [21], MiniGPT-v2 [3], Ferret [44], Kosmos-2 [31], Griffon [48], Qwen-VL [1], CogVLM [41], and LLaVA [24]. Detailed model specifications are provided in the Supplementary Material.
Zero-shot evaluation details. For zero-shot evaluation, we use the original model checkpoints as provided in their respective papers. The implementations and checkpoints for GLIP and GDINO are obtained from the MMDetection [4].
Fine-tuning evaluation details. Our fine-tuning protocol maintains consistency across all experiments to ensure a fair comparison. We preserve the original learning strategies while excluding random crop augmentation due to its negative effect on position-sensitive samples. For LMMs, we employ the LoRA [11] fine-tuning strategy and follow the instruction tuning data structures. All fine-tuning experiments run for 5 epochs using 4 NVIDIA A100 GPUs.
5.3 Experimental results
Table 2: Experimental results of zero-shot baselines on RefDrone benchmark. Top groups: specialist models. Bottom groups: LMMs for REC tasks. GDINO-T and GDINO-B denote GroundingDINO [28] with Swin-Tiny and Swin-Base backbones. The best results in each group are denoted with bold.
Methods F1inst. Accinst. F1img. Accimg. MDETR [14] 8.42 4.41 2.99 1.63 GLIP [19] 5.46 3.84 9.20 8.54 GDINO-T [28] 1.18 1.84 3.94 6.35 GDINO-B [28] 1.97 2.23 6.43 7.58 Shikra [5] 0.80 0.52 2.26 1.60 ONE-PEACE [38] 1.02 0.51 2.64 1.34 SPHINX-v2 [21] 1.59 0.80 4.61 2.36 MiniGPT-v2 [3] 2.69 1.36 6.38 3.29 Ferret [44] 3.18 1.62 8.48 4.43 Kosmos-2 [31] 8.06 4.20 8.64 4.52 Griffon [48] 9.16 4.81 16.77 9.18 Qwen-VL [1] 10.91 5.77 18.32 10.08 CogVLM [41] 15.38 8.33 30.73 18.15
Zero-shot results. Table 2 presents the zero-shot evaluation results, assessing the models’ domain generalization capability. CogVLM [41] demonstrates state-of-the-art performance across multiple metrics. However, several advanced models, including Shikra [5], ONE-PEACE [38], and SPHINX-v2 [21], are limited to outputting only a single bounding box, due to constraints in their pre-training data or output strategy. This restriction significantly impacts their performance in multi-target scenarios.
Table 3: Experimental results of fine-tuning baselines on RefDrone benchmark. Top groups: specialist models. Bottom groups: LMMs for REC tasks.
Methods F1inst. Accinst. F1img. Accimg. MDETR [14] 32.60 19.49 19.17 10.81 GLIP [19] 24.23 14.86 16.92 13.29 GDINO-T [28] 30.50 19.02 29.65 21.07 NGDINO-T (Ours) 33.34 20.98 32.45 22.84 GDINO-B [28] 31.96 19.99 31.69 22.26 NGDINO-B (Ours) 34.44 21.76 34.01 23.89 MiniGPT-v2 [3] 4.97 2.74 13.56 8.97 LLaVA [24] 6.00 3.63 14.43 11.57 Qwen-VL [1] 14.14 7.61 20.10 11.17 RDAgent (Ours) 58.14 41.13 37.07 23.54
Fine-tuning results. Table 3 illustrates the fine-tuning performance across different methods. The specialist model MDETR [14] shows strong performance in instance-level metrics, achieving comparable results to GDINO-B [28], but struggles in image-level understanding. Conversely, LMMs like Qwen-VL [1] demonstrate superior image-level comprehension (F1img.: 20.10%, Accimg.: 11.17%) while struggling with instance-level tasks (F1inst.: 14.14%, Accinst.: 7.61%). This performance disparity can be attributed to LMMs’ strong global image understanding capabilities but limited effectiveness in detecting small objects due to input resolution constraints.
The proposed RDAgent achieves superior results compared to existing approaches, particularly in instance-level metrics, surpassing GDINO-B [28] by 26.18% in F1inst. and 21.14% in Accinst.. The relatively smaller gains in image-level metrics (5.38% in F1img. and 1.28% in Accimg.) highlight the importance of instance-level metrics. However, RDAgent’s multi-step pipeline and reliance on ChatGPT responses result in longer processing time compared to end-to-end approaches.
The proposed end-to-end method, NGDINO, consistently improves over the baseline GDINO [28] across both backbone architectures. Figure 7 shows the visualization results comparing NGDINO and baseline GDINO, demonstrating NGDINO’s effectiveness with multi-target samples.
Table 4: Experimental results on gRefCOCO [22] dataset. Asterisk (*) denotes results reported in the gRefCOCO paper.
Methods testA testB
Pr@0.5 $\uparrow$ N-acc. $\uparrow$ Pr@0.5 $\uparrow$ N-acc. $\uparrow$
MDETR* [14] 50.0 34.5 36.5 31.0
UNINEXT* [43] 46.4 49.3 42.9 48.2
GDINO-T [28] 45.69 79.02 44.83 76.69
NGDINO-T 46.05 83.17 45.57 78.08
To further validate the effectiveness of the proposed NGDINO, we conduct additional experiments on the gRefCOCO [22] dataset, which also includes multi-target and no-target samples (Table 4). NGDINO-T outperforms the baseline method GDINO-T, particularly in N-acc. metrics (by 4.15% in test A and 1.39% in test B), which evaluate accuracy on no-target samples. The limited improvements in Pr@0.5 metrics can be attributed to the relatively simple nature of multi-target samples in gRefCOCO, which typically contain only two objects per sample.
Figure 7: Visualization results comparing NGDINO and GDINO [28], where red boxes indicate true positives and yellow boxes denote false negatives.
Figure 8: Failure cases of NGDINO on RefDrone dataset. Red, green, and yellow boxes indicate true positives, false positives, and false negatives, respectively.
Table 5: Ablation study on RefDrone dataset for the effectiveness of each component in NGDINO, with Swin-Tiny backbone. NPH denotes number prediction head. NCA denotes number cross-attention.
NPH NCA Acc_$\_inst.$ Acc_$\_img.$ FPS 19.02 21.07 13.5 19.09 (0.07$\uparrow$) 21.18 (0.11$\uparrow$) 12.8 (0.7$\downarrow$) 19.51 (0.49$\uparrow$) 21.71 (0.53$\uparrow$) 12.9 (0.6$\downarrow$) 20.98 (1.96$\uparrow$) 22.84 (1.77$\uparrow$) 12.3 (1.2$\downarrow$)
Table 6: Ablation study for the impact of varying selected number query length. Params indicates additional parameters introduced.
Length F1inst. Accinst. F1img. Accimg. Params 1 31.63 19.80 31.23 22.12 1.58M 10 33.34 20.98 32.45 22.84 1.65M 100 32.06 20.10 32.23 22.81 2.34M
Table 7: Ablation study for the effectiveness of RDAgent.
Methods F1inst. Accinst. F1img. Accimg. ReCLIP [36] 24.62 14.04 11.58 6.15 GPT4-o 52.38 35.65 35.50 22.38 RDAgent 58.14 41.13 37.07 23.54
5.4 Ablation studies
Ablations of NGDINO components. Table 5 presents the analysis of each component in NGDINO. The number prediction head alone shows a slight increase in Accinst. from 19.02% to 19.09%, indicating minimal impact of this auxiliary head on the referring tasks. When introducing the number cross-attention without number selection, we observe a more substantial improvement, with Accinst. increasing from 19.02% to 19.51%. This improvement is attributed to the additional parameters introduced in the decoder. The significant improvement is achieved through the combination of both the number prediction task and number cross-attention components. This contribution increases Accinst. from 19.02% to 20.98%. While the additional computational cost results in a minor decrease in inference speed (FPS), this trade-off yields a 1.96% increase in Accinst..
Ablations of query length. Table 6 analyzes the impact of varying the selected number query length. Utilizing a minimal query length of 1 lacks the capacity to capture complex numerical information. Conversely, extending the query length to 100 increases parameter count and introduces optimization difficulties, thereby adversely affecting performance. Through these experiments, we determine that a query length of 10 provides an optimal trade-off.
Number prediction head performance. The number prediction head achieves an overall number prediction accuracy of 53.5%. During the bounding box prediction stage, this accuracy is observed to be 22.84%. This performance gap demonstrates the effectiveness of number prediction head. Besides, the number prediction achieves a mean absolute error (MAE) of 0.51, suggesting high precision as predictions closely align with ground truth values.
Ablations of RDAgent. Table 7 analyzes the effectiveness of the proposed RDAgent. The GPT4-o baseline is established using only Step 3 of RDAgent. We also compare against ReCLIP [36], a two-stage instance ranking method using CLIP [33] for REC tasks. All experiments utilize Faster-RCNN [34] for object detection. Results show that RDAgent consistently outperforms GPT4-o and ReCLIP across all metrics. However, the performance is partially limited by the Faster-RCNN object detector, which achieves only 18.0 mAP on the VisDrone2019-DET [49] dataset.
5.5 Limitations
While NGDINO addresses the multi-target and no-target challenges, several challenges remain. Figure 8 shows typical NGDINO failure cases, highlighting challenges in the RefDrone dataset. The examples show rich contextual expressions, challenging backgrounds, and small-scale object detection. These challenging cases reflect real-world applications and highlight the areas for future improvement.
6 Conclusion
In this work, we introduce RefDrone, a challenging benchmark specifically designed for referring expression comprehension in drone scenes. The dataset is constructed using the proposed RDAgent, a semi-auto annotation framework leveraging a multi-agent system. Furthermore, we develop NGDINO to address the multi-target and no-target challenges in the RefDrone dataset. In the future, we aim to further enhance NGDINO to address additional challenges presented by RefDrone. We also plan to expand the benchmark to include referring expression segmentation and referring expression tracking tasks.
Supplementary Material
We provide supplementary material for further analysis, organized as follows:
  1. 1.
    Color attribute analysis of RefDrone. (Appendix A)
  2. 2.
    Details of baseline method. (Appendix B)
  3. 3.
    Additional dataset examples. (Appendix C)
  4. 4.
    Resutlts w.r.t different object scales. (Appendix D)
  5. 5.
    Prompts and examples in RDAgent. (Appendix E)
Appendix A Dataset color attribute analysis
In the RefDrone dataset, color attributes are present in 69% expressions. We initially use the HSV color space to define six primary colors: white, black, red, blue, green, and yellow. During the annotation process, RDAgent expands the color set by adding six more colors: orange, pink, grey, purple, brown, and silver. The distribution of these color terms across the expressions is illustrated in Figure 9.
Figure 9: Distribution of color terms in RefDrone expressions.
Appendix B Details of baseline methods
The details for each baseline method:
  • Specialist models:
    • MDETR: ResNet-101 with BERT-Base,
      pretrained on Flickr30k, RefCOCO/+/g, VG.
    • GLIP: Swin-Tiny with BERT-Base,
      pretrained on Object365.
    • GDINO-T: Swin-Tiny with BERT-Base,
      pretrained on Object365, GoldG, GRIT, V3Det.
    • GDINO-B: Swin-Base with BERT-Base,
      pretrained on Object365, GoldG, V3Det.
  • Large multimodal models:
    • Kosmos-2: 1.6B parameters.
    • ONE-PEACE: 4B parameters, visual grounding API.
    • Shikra: 7B parameters, delta-v1 version.
    • MiniGPT-v2: 7B parameters.
    • LLaVA: 7B parameters, v1.5 version.
    • Qwen-VL: 7B parameters.
    • Ferret: 7B parameters.
    • CogVLM: 7B parameters, grounding-specific version.
    • SPHINX-v2, Griffon: 13B parameters.
Appendix C Dataset examples
To provide a comprehensive understanding of our RefDrone dataset, we present representative examples in Figure 10. These samples demonstrate the three key challenges in our dataset, highlighting its real-world applicability.
Appendix D The results of different object scales.
We add small/medium/large instance accuracy metrics (ACCs, ACCm, ACCl) of RefDrone dataset in Table 8.
Table 8: Fine-tuned results w.r.t. different object scales.
Methods Accs Accm Accl Methods Accs Accm Accl GLIP 2.65 18.28 27.29 LLaVA 4.18 7.26 11.99 GDINO-T 8.56 20.66 33.99 Qwen-VL 0.73 10.60 30.18 NGDINO-T 10.84 23.38 40.13 RDAgent 32.54 47.97 47.17
Appendix E Prompts and examples for RDAgent
In this section, we provide the prompts and examples employed in RDAgent. Table 9 presents the prompt construction process for expression generation (Step 3), which includes the system prompt and few-shot in-context learning examples. One in-context learning example is illustrated in Table 10. The system prompts used for each step are detailed in Table 11. Additionally, the system prompts for the feedback mechanism are presented in Table 12.
Figure 10: Dataset samples from RefDrone.
Table 9: Illustration of RDAgent’s prompt construction for expression generation (Step 3). Few-shot in-context-learning examples are from fewshot_samples. A representative example is shown in Figure 10
messages = [ {"role":"system", "content": f"""As an AI visual assistant, your role involves analyzing a single image. You are supplied with three sentences that caption the image, along with additional data about specific attributes of objects within the image. This can include information about categories, colors, and precise coordinates. Such coordinates, represented as floating-point numbers that range from 0 to 1, are shared as center points, denoted as (x, y), identifying the center x and y. When coordinate x tends to 0, the object nears the left side of the image, shifting towards the right as coordinate x approaches 1. When coordinate y tends to 0, the object nears the top of the image, shifting towards the bottom as coordinate y approaches 1. Your task is to classify the provided objects based on various characteristics, while also substantiating your classification. This classification should be thoroughly justified, with criteria including but not limited to relationships or relative locations of objects. To refer to a specific object, use the provided coordinates directly. Base your classification justifications on direct observations from the image, avoiding any hypothesizing or assumptions."""} ] for  sample in fewshot_samples:
       messages.append({"role":"user", "content":sample[‘context’]})
messages.append({"role":"assistant", "content":sample[‘response’]} )
[…]
Table 10: One example to illustrate the input and output in RDAgent Step 3 (expression generation). The top block shows the input contexts such as image, object, caption and color attributed. The bottom block shows the generated annotations.
Context type 1: Captions
The image shows an aerial view of a busy urban street with multiple lanes of traffic, flanked by parking areas and buildings. A median strip with greenery runs down the center, dividing the opposite flows of vehicles. On either side of the road, there are commercial and residential structures. Parking lanes on both sides of the main road are filled with parked vehicles.
[…]
Response : The blue cars park on the right parking lot. Coordinates: [0.628, 0.478], [0.646, 0.692], [0.649, 0.727] The white cars park on both sides of the main road in a line. Coordinates: [0.431, 0.051], [0.431, 0.15], [0.431, 0.101], [0.434, 0.256], [0.439, 0.523], [0.443, 0.813], [0.445, 0.902], [0.446, 0.982], [0.235, 0.43], [0.232, 0.486], [0.223, 0.674], [0.215, 0.904]
Table 11: System prompts employed at different stages of RDAgent, replacing the default system messages shown in Table 9
System prompt: Step 1 scene understanding
You are an AI visual assistant that specializes in providing clear and accurate descriptions of images without any ambiguity or uncertainty. Your descriptions should focus solely on the content of the image itself and avoid mentioning any location-specific details such as regions or countries where the image might have been captured.
System prompt: Step 2 color categorization
As an AI visual assistant, your role involves analyzing a single image.
You are supplied with the specific attributes of objects within the image. This can include information about categories, colors, and precise coordinates. Such coordinates, represented as floating-point numbers that range from 0 to 1, are shared as center points, denoted as (x, y), identifying the center x and y. Your task is to assess whether the given colors of specific objects match their appearance in the image. Respond with ”Yes” when the colors are appropriate. In cases where the colors are deemed inappropriate, respond with a concise ”No.” System prompt: Step 3 expression generation
As an AI visual assistant, your role involves analyzing a single image. You are supplied with three sentences that caption the image, along with additional data about specific attributes of objects within the image. This can include information about categories, colors, and precise coordinates. Such coordinates, represented as floating-point numbers that range from 0 to 1, are shared as center points, denoted as (x, y), identifying the center x and y. When coordinate x tends to 0, the object nears the left side of the image, shifting towards the right as coordinate x approaches 1. When coordinate y tends to 0, the object nears the top of the image, shifting towards the bottom as coordinate y approaches 1.
Your task is to classify the provided objects based on various characteristics, while also substantiating your classification. This classification should be thoroughly justified, with criteria including but not limited to relationships or relative locations of objects. To refer to a specific object, use the provided coordinates directly. Base your classification justifications on direct observations from the image, avoiding any hypothesizing or assumptions. System prompt: Step 4 quality evaluation
As an AI visual assistant, your role involves analyzing a single image. You are supplied with three sentences that describe the image, along with additional data about specific attributes of objects within the image. This can include information about categories, colors, and precise coordinates. Such coordinates, represented as floating-point numbers that range from 0 to 1, are shared as center points, denoted as (x, y), identifying the center x and y. When coordinate x tends to 0, the object nears the left side of the image, shifting towards the right as coordinate x approaches 1. When coordinate y tends to 0, the object nears the top of the image, shifting towards the bottom as coordinate y approaches 1. Besides, you are supplied with the description of the objects and their corresponding attributes.
Your task is to confirm whether the description exclusively relates to the described objects without including any others in the visual. Respond ”yes” if it matches, or ”no” with an explanation if it does not.
Table 12: RDAgent system prompts for feedback mechanism. Differences from Table 11 are highlighted
System prompt: Step 2 color categorization with feedback mechanism
As an AI visual assistant, your role involves analyzing a single image.
You are supplied with the specific attributes of objects within the image. This can include information about categories, colors, and precise coordinates. Such coordinates, represented as floating-point numbers that range from 0 to 1, are shared as center points, denoted as (x, y), identifying the center x and y. Your task is to assess whether the given colors of specific objects match their appearance in the image. Respond with ”Yes” when the colors are appropriate. In cases where the colors are deemed inappropriate, respond with a concise ”No.” System prompt: Step 3 expression generation with feedback mechanism
As an AI visual assistant, your role involves analyzing a single image. You are supplied with three sentences that caption the image, along with additional data about specific attributes of objects within the image. This can include information about categories, colors, and precise coordinates. Such coordinates, represented as floating-point numbers that range from 0 to 1, are shared as center points, denoted as (x, y), identifying the center x and y. When coordinate x tends to 0, the object nears the left side of the image, shifting towards the right as coordinate x approaches 1. When coordinate y tends to 0, the object nears the top of the image, shifting towards the bottom as coordinate y approaches 1. You are also provided with descriptions and the objects that initially failed to match, along with the reasons for the discrepancies.
Your task is to revise both the description and the corresponding objects to correct these mismatches based on the provided reasons. Ensure that the revised description accurately matches the corresponding objects depicted in the visual content.
System prompt: Step 4 quality evaluation with feedback mechanism
As an AI visual assistant, your role involves analyzing a single image. You are supplied with three sentences that describe the image, along with additional data about specific attributes of objects within the image. This can include information about categories, colors, and precise coordinates. Such coordinates, represented as floating-point numbers that range from 0 to 1, are shared as center points, denoted as (x, y), identifying the center x and y. When coordinate x tends to 0, the object nears the left side of the image, shifting towards the right as coordinate x approaches 1. When coordinate y tends to 0, the object nears the top of the image, shifting towards the bottom as coordinate y approaches 1. Besides, you are supplied with the description and the objects that initially failed to match.
Your task is to provide detailed reasoning for unsuccessful object matches.

### v2

@@
 Abstract
-Drones have become prevalent robotic platforms with diverse applications, showing significant potential in Embodied Artificial Intelligence (Embodied AI).
-Referring Expression Comprehension (REC) enables drones to locate objects based on natural language expressions, a crucial capability for Embodied AI.
-Despite advances in REC for ground-level scenes, aerial views introduce unique challenges including varying viewpoints, occlusions and scale variations.
-To address this gap, we introduce RefDrone, a REC benchmark for drone scenes.
-RefDrone reveals three key challenges in REC: 1) multi-scale and small-scale target detection; 2) multi-target and no-target samples; 3) complex environment with rich contextual expressions.
-To efficiently construct this dataset, we develop RDAgent (referring drone annotation framework with multi-agent system), a semi-automated annotation tool for REC tasks.
-RDAgent ensures high-quality contextual expressions and reduces annotation cost.
-Furthermore, we propose Number GroundingDINO (NGDINO), a novel method designed to handle multi-target and no-target cases.
-NGDINO explicitly learns and utilizes the number of objects referred to in the expression.
-Comprehensive experiments with state-of-the-art REC methods demonstrate that NGDINO achieves superior performance on both the proposed RefDrone and the existing gRefCOCO datasets.
-The dataset and code will be publicly at https://github.com/sunzc-sunny/refdrone.
+Drones have become prevalent robotic platforms with significant potential in Embodied AI.
+A crucial capability for drone-based Embodied AI is Referring Expression Comprehension (REC), which enables locating objects with language expressions.
+Despite advances in REC for ground-level scenes, drones’ unique capability for broad observation introduces distinct challenges: multiple potential targets, small-scale objects, and complex environmental contexts.
+To address these challenges, we introduce RefDrone, an REC benchmark for drone scenes.
+RefDrone reveals three key challenges: 1) multi-target and no-target scenarios; 2) multi-scale and small-scale target detection; 3) complex environment with rich contextual reasoning.
+To efficiently construct this dataset, we develop RDAnnotator, a semi-automated annotation framework where specialized modules and human annotator collaborate through feedback loops.
+RDAnnotator ensures high-quality contextual expressions while reducing annotation costs.
+Furthermore, we propose Number GroundingDINO (NGDINO), a novel method to handle multi-target and no-target cases.
+NGDINO explicitly estimates the number of objects referred to in the expression and incorporates this numerical pattern into the detection process.
+Comprehensive experiments with state-of-the-art REC methods demonstrate that NGDINO achieves superior performance on RefDrone, as well as on the general-domain gRefCOCO and remote sensing RSVG benchmarks.
+The data and code are completely available at https://github.com/sunzc-sunny/refdrone.
 1 Introduction
 Figure 1: Examples of the various challenges in RefDrone dataset.
-Table 1: Comparison of referring expression comprehension datasets: RefCOCO/+/g [45, 29], gRefCOCO [22], D3 [42], RIS-CQ [13], RSVG [47], and RefDrone (ours).
-Avg.
-object: average number of objects per expression.
-Avg.
-length: words per expression.
-No target: expressions without referred objects.
-Expression type: expression generation method.
-Small target: small-scale target detection challenge.
-RefCOCO/+/g [45, 29] gRefCOCO [22] D3 [42] RIS-CQ [13] RSVG [47] RefDrone Image source COCO [20] COCO [20] COCO [20] VG [16]+COCO [20] DIOR [18] VisDrone [49] Avg.
-object 1.0 1.4 1.3 3.6 2.2 3.8 Avg.
-length 3.6/3.5/8.4 4.9 6.3 13.2 7.5 9.0 No target Expression type Manual Manual Manual LLM (GPT3.5-t) Templated LMM (GPT4-o) Small target
-Drones/UAVs have become increasingly popular in our daily lives, serving both personal and professional purposes, such as entertainment, package delivery, traffic surveillance, and emergency rescue.
-Their ability to move freely and their broad observation capabilities make them important platforms for Embodied AI [6, 27, 17, 8, 25].
+Drones/UAVs have become increasingly popular in our daily lives, serving both personal and professional purposes, such as entertainment, package delivery, traffic surveillance, and emergency rescue [1, 42, 45].
+Their ability to move freely and to observe broadly make them important platforms for Embodied AI applications [12, 35, 23, 14, 31].
 A crucial capability in Embodied AI is Referring Expression Comprehension (REC) [43, 48, 4, 49], which serves as a critical bridge between natural language understanding and visual perception.
 REC requires drones to localize specific objects in images based on natural language expressions.
-However, existing REC datasets primarily focus on ground-level perspectives, such as the RefCOCO [45, 29] dataset.
-The application of REC in drone-based scenarios presents unique challenges, including extreme viewpoint variations, occlusions, and scale variations across objects.
-In this work, we introduce RefDrone, a challenging benchmark designed for REC in drone scenes.
-The RefDrone dataset contains 17,900 referring expressions across 8,536 images, comprising 63,679 object instances.
-As illustrated in Figure 1, RefDrone poses three key challenges that distinguish it from existing datasets: (1) multi-scale and small-scale target detection, with 31% small objects and 14% large objects; (2) multi-target and no-target samples, where expressions can refer to any number of objects from 0 to 242; and (3) complex environment with rich contextual expressions.
-The environmental complexity manifests in varied viewpoints, diverse lighting conditions, and complex backgrounds.
-The contextual richness is reflected in expressions that describe spatial relations, attributes, and interactive relationships.
-To efficiently construct this comprehensive dataset, we develop RDAgent (referring drone annotation framework with multi-agent system), a semi-automated annotation pipeline.
-RDAgent restructures traditional annotation workflows into an interactive system of multiple agents and human annotators.
-In this framework, agents with diverse roles collaborate through feedback loops to generate and verify annotations, while human involvement is minimized to quality control and minor adjustments.
-RDAgent significantly reduces annotation costs while maintaining high-quality standards for complex expressions, and it can be extended to generate annotations for other REC tasks.
-Furthermore, we propose a novel method called Number GroundingDINO (NGDINO).
-Our key insight is that utilizing the number information of referred objects enhances the handling of multi-target and no-target samples.
-NGDINO comprises three key components: (1) a number prediction head that estimates target object quantities, (2) a set of learnable number-queries encoding numerical patterns across different quantity samples, and (3) a number cross-attention module that integrates number queries with detection queries.
-Extensive experiments on both our drone-specific RefDrone benchmark and the general-domain gRefCOCO [22] dataset reveal that leveraging the number information greatly improves performance on multi-target and no-target samples.
+However, existing REC datasets primarily focus on ground-level perspectives, such as RefCOCO [54, 36] dataset.
+The application of REC in drone-based scenarios presents unique challenges.
+In this work, we introduce RefDrone, a challenging REC benchmark for drone scenes.
+RefDrone comprises 17,900 referring expressions annotated over 8,536 images, with 63,679 object instances.
+As illustrated in Figure 1, RefDrone poses three primary challenges that distinguish it from existing datasets: (1) multi-target and no-target samples, where expressions may refer to any number of objects, ranging from 0 to 242; (2) multi-scale and small-scale target detection, with 31% small objects and 14% large objects; and (3) complex environment with rich contextual reasoning, encompassing diverse viewpoints, lighting conditions, and intricate backgrounds.
+The referring expressions richly capture spatial relations, object attributes, and inter-object interactions.
+As summarized in Table 1, RefDrone offers greater diversity and complexity than previous REC benchmarks.
+We further evaluate 13 representative REC models using both fine-tuning and zero-shot settings.
+Notably, all models exhibit significantly poorer performance on RefDrone compared to standard REC datasets (e.g., GroundingDINO [34] yields 21.07 $Acc_{img.}$ on RefDrone vs.
+91.4 $Acc_{img.}$ on RefCOCO), highlighting the inherent difficulty and unique challenges posed by our benchmark.
+To efficiently construct RefDrone, we develop RDAnnotator, a semi-automated annotation pipeline for referring expression annotation in drone scenes.
+RDAnnotator leverages multiple specialized LMM-based modules, which collaborate within a feedback loop to generate and validate annotations.
+This tool reduces human involvement to quality control and minor adjustments, thereby significantly lowering annotation costs, while maintaining high-quality standards for complex expressions.
+RDAnnotator achieves a cost of just \$0.0539 per expression (GPT-4o API usage) and reduces human annotation time to under one minute per expression.
+This cost-efficiency makes RDAnnotator a scalable solution for large-scale dataset construction and can be readily adapted to other REC tasks.
+Furthermore, we propose Number GroundingDINO (NGDINO) to address the challenges of multi-target and no-target referring expressions.
+Our key insight is that explicitly utilizing the number information of referred objects enhances the handling of these scenarios.
+NGDINO includes three key components: (1) a number prediction head that estimates the count of target objects, (2) learnable number-queries that capture numerical patterns corresponding to varying object quantities, and (3) a number cross-attention module that fuses number queries with detection queries to inform object localization.
+Extensive experiments on both our RefDrone benchmark, as well as the general-domain gRefCOCO [28] and remote sensing RSVG [56] datasets, demonstrate that NGDINO substantially improves performance, particularly in the challenging multi-target and no-target cases.
 In summary, our contributions are listed as follows:
-  1. 1.
-    RefDrone Benchmark: The first comprehensive benchmark for referring expression comprehension in drone scenes.
-RefDrone poses three key challenges and provides comprehensive baseline evaluations across both specialized models and large multimodal models.
-  2. 2.
-    RDAgent Annotation Framework: A novel semi-automated annotation framework employs a multi-agent system. RDAgent reduces annotation costs while ensuring high-quality referring expressions.
-  3. 3.
-    NGDINO Method: A novel method specifically designed to handle multi-target and no-target samples through three components: a number prediction head, learnable number-queries, and number cross-attention.
-NGDINO achieves superior performance on both RefDrone and gRefCOCO datasets.
+  • RefDrone Benchmark: We introduce RefDrone, a comprehensive benchmark for referring expression comprehension in drone scenes.
+RefDrone poses three key challenges: multi-target/no-target scenarios, multi-scale/small-object detection, and complex contextual reasoning.
+We provide thorough baseline evaluations with 13 representative REC models.
+  • RDAnnotator Framework: We propose RDAnnotator, a semi-automated annotation framework that significantly reduces human effort while ensuring high-quality referring expressions.
+This framework is scalable for large-scale dataset construction and generalizable beyond drone imagery.
+  • NGDINO Method: We develop NGDINO, a novel approach that explicitly models the number of referred objects to address multi-target and no-target cases.
+NGDINO achieves state-of-the-art performance on RefDrone and shows consistent improvements on gRefCOCO and RSVG benchmarks.
+Table 1: Comparison of REC datasets relevant to RefDrone.
+Avg.
+objects: mean objects per expression.
+Avg.
+length: mean words per expression.
+No target: expressions without referred objects.
+Expression type: method for expression generation.
+Small target: percentage of small-scale objects.
+RefCOCO/+/g [54, 36] gRefCOCO [28] D3 [51] RIS-CQ [19] RSVG [56] RefDrone (Ours) Image source COCO [26] COCO [26] COCO [26] VG [22]+COCO [26] DIOR [24] VisDrone [60] Avg.
+objects 1.0 1.4 1.3 3.6 2.2 3.8 Avg.
+length 3.6/3.5/8.4 4.9 6.3 13.2 7.5 9.0 No target Expression type Manual Manual Manual LLM (GPT3.5-t) Templated LMM (GPT4-o) Small target 0/0/0% 0.1% 6.3% - 17.2% 31.1%
 2 Related Works
 2.1 Referring expression understanding datasets
-Referring expression understanding identifies specific regions within images or videos using natural language expressions.
-Two primary subtasks are Referring Expression Comprehension (REC) and Referring Expression Segmentation (RES).
-REC outputs detection bounding boxes, while RES outputs segmentation masks.
-Early datasets such as ReferIt [15] and RefCOCO [45] are pioneering but are limited to single-target expressions and simple expressions.
-Subsequently, datasets [22, 37, 13, 42] introduce more challenges.
+Referring expression understanding identifies regions in visual content from natural language expressions.
+The main subtasks are Referring Expression Comprehension (REC), which outputs detection bounding boxes, and Referring Expression Segmentation (RES), which outputs segmentation masks.
+Datasets in this domain show increasing visual and linguistic complexity.
+Early datasets such as ReferIt [21] and RefCOCO [54] are pioneering but limited to simple expressions and single-target scenarios.
+Later datasets [32, 9, 28, 46, 19, 51] introduce more challenges.
 gRefCOCO [28] supports multi-target expressions.
-OV-VG [37] enables an open-vocabulary setting.
-D3 [42] and RIS-CQ [13] introduce more complex expressions.
-Additionally, domain-specific datasets [47, 30, 32] emerge to address unique applications.
-RSVG [47] focuses on remote sensing, RAVAR [30] targets video action recognition, and RIO [32] addresses affordance detection.
-However, drone scenes remain unexplored in REC tasks, which motivates our introduction of the RefDrone benchmark.
-Recent advances in dataset creation have been significantly influenced by large language models (LLMs).
-Notable works such as LLaVA [24] and Ferret [44] show the effectiveness of LLMs in generating instruction tuning data.
-Similarly, RIS-CQ [13] and RIO [32] employ LLMs to generate complex referring expressions.
-However, these approaches have two limitations.
-First, LLMs are used mainly as text generators, potentially producing expressions without proper grounding in visual context.
-Second, the lack of iterative feedback mechanisms in the annotation pipelines may degrade annotation quality.
-Figure 2: The overview of the RefDrone annotation process with RDAgent.
-Multiple specialized agents collaborate both with each other and human annotators through iterative feedback loops to generate high-quality annotations.
+OV-VG [46] enables open-vocabulary queries.
+D3 [51] and RIS-CQ [19] offer more complex expressions.
+Domain-specific datasets [10, 8, 56, 37, 39] address specialized applications: RSVG [56] for remote sensing, RAVAR [37] for video action recognition, RIO [39] for affordance detection.
+However, drone scenes remain unexplored in REC, motivating our RefDrone benchmark.
+Table 1 compares RefDrone with other related REC benchmarks.
+Recent dataset creation approaches leverage large language models (LLMs) to reduce annotation costs.
+Works such as LLaVA [30] and Ferret [53] demonstrate the effectiveness of LLMs in generating instruction tuning data.
+Similarly, RIS-CQ [19] and RIO [39] employ LLMs to generate referring expressions.
+However, these approaches treat LLMs as text generators and may lack proper visual grounding.
+In contrast, we employ a semi-automated annotation framework with iterative feedback mechanisms throughout the annotation pipeline to reduce annotation costs and improve quality.
 2.2 Referring expression comprehension methods
-Referring expression comprehension methods can be broadly categorized into specialist models and large multimodal models (LMMs).
-Specialist models include two-stage and one-stage methods.
-Two-stage methods [12, 23, 10, 36, 9] typically approach REC as a ranking task, first generating region proposals and then ranking them based on language input.
-However, they are often criticized for slow inference speeds.
-In contrast, one-stage methods [14, 19, 26, 7, 43, 28] directly predict target regions guided by language.
-These approaches leverage transformer architectures to enable cross-modal interactions between visual and textual features.
-Among these methods, GroundingDINO (GDINO) [28] has gained widespread attention for its impressive results in both open-set object detection and REC tasks.
-However, GDINO lacks explicit mechanisms to handle multi-target or no-target scenarios.
-Recent studies have increasingly evaluated LMMs [24, 44, 3, 5, 1, 31, 38, 21, 48, 41] on REC tasks to assess their visual-language understanding capabilities.
-These models leverage large-scale referring instruction tuning data to achieve remarkable performance.
+REC methods can be broadly categorized into large multimodal models (LMMs) and specialist models.
+Large Multimodal Models [30, 53, 5, 7, 3, 38, 47, 27, 57, 50] have recently been applied to REC tasks as part of evaluating their broader visual-language understanding capabilities.
+These models leverage extensive referring instruction tuning data to achieve competitive performance without task-specific architectural designs.
 Despite their broad capabilities, LMMs consistently struggle with small object detection, primarily due to their inherent input resolution constraints.
 This constraint forces image downsampling, resulting in the loss of fine-grained visual details.
+Specialist models include two-stage and one-stage methods.
+Two-stage methods [18, 29, 16, 44, 15] typically approach REC as a ranking task: first generating region proposals through a detector, then ranking them based on their alignment with the language input.
+However, they are often criticized for slow inference speeds.
+In contrast, one-stage methods [20, 25, 33, 13, 52, 34] directly predict target regions guided by language input.
+These approaches leverage transformer to enable cross-modal interactions between visual and textual features.
+Among these methods, GroundingDINO (GDINO) [34] has gained widespread attention for its impressive results in REC tasks.
+Our work extends GDINO by incorporating explicit number modeling to handle multi-target and no-target scenarios, which are critical challenges in drone-based REC tasks.
 3 RefDrone benchmark
+Figure 2: The overview of the RefDrone annotation process with RDAnnotator.
+Multiple specialized LMM-based modules collaborate both with each other and human annotators through iterative feedback loops to generate high-quality annotations.
 3.1 Data source
-The RefDrone dataset builds upon VisDrone2019-DET [49], a drone-captured dataset for object detection.
-The source images are collected across multiple scenarios, illumination conditions, and flying altitudes.
-To ensure meaningful visual content, we implement two filtering criteria.
-At the image level, we exclude images containing less than 3 objects.
-At the object level, we exclude objects with bounding box areas below 64 pixels.
-The original VisDrone2019-DET annotations provide object categories and bounding box coordinates.
-We convert the bounding box to a normalized center point (range 0-1).
+The RefDrone benchmark is built upon VisDrone2019-DET [60], a large-scale high-quality drone-captured object detection dataset.
+Source images are collected across multiple scenarios, illumination conditions, and flying altitudes.
+To ensure meaningful visual content, we filter out images with fewer than three objects and objects with bounding box areas below 64 pixels.
+VisDrone2019-DET provides object categories and bounding box coordinates, which we convert to normalized center points (range 0-1).
 This approach reduces the input context for LMMs while preserving spatial information.
-3.2 RDAgent for semi-automated annotation
-RDAgent (referring drone annotation framework with multi-agent system) is a semi-automated annotation framework that enhances traditional annotation pipelines by integrating LMMs with human annotators.
-Central to RDAgent are multiple LMM agents, specifically GPT-4o, each perform specialized tasks via distinct system prompts.
-As shown in Figure 2, RDAgent employs five structured steps.
+3.2 RDAnnotator for semi-automated annotation
+RDAnnotator (referring drone annotator) is a semi-automated annotation framework that minimizes human effort while ensuring high quality.
+It integrates specialized modules powered by LMMs with human validation.
+As shown in Figure 2, RDAnnotator employs five structured steps.
 Step 1: scene understanding.
-A GPT-4o-based captioning agent generates three diverse textual descriptions per image.
-These captions provide contextual perspectives (e.g., object relationships, spatial layouts) to anchor subsequent referring expression generation.
-Figure 3: Object size distribution analysis.
-(a) Object size distribution in RefDrone dataset (small: $<32^{2}=1024$ pixels, normal: 1024 to 9216 pixels, large: $>96^{2}=9216$ pixels).
-(b) Object size histograms in RefDrone and gRefCOCO [22] datasets.
+A GPT-4o-based scene parsing module generates three diverse textual descriptions per image.
+These captions establish spatial layouts, object relationships, and contextual foundations to guide subsequent referring expression generation.
 Step 2: color categorization.
-Color attributes serve as crucial discriminative features in referring expressions.
-Color attributes are extracted using a hybrid pipeline: (1) a WideResNet-101 [46] model trained on HSV color space-based labels, refined via manual validation, and (2) a dedicated LMM agent that verifies predictions to mitigate illumination- or occlusion-induced errors.
+Color attributes are crucial discriminative features in referring expressions.
+We implement a hybrid color extraction pipeline, which combines low-level CNN-based classification (WideResNet-101 classifier [55]) with high-level LMM semantic reasoning, ensuring robust color attribution in challenging aerial conditions.
 Step 3: expression generation.
-To address the complexity in referring expression generation, we reformulate the referring expression generation task as an object grouping problem.
-The target of the agent is to group semantically related objects and to provide appropriate reasons for each group.
+To address the complexity of referring expression generation, we reformulate the referring expression generation task as an object grouping problem.
+The target of the module is to group semantically related objects and provide appropriate reasons for each group.
 The reasons serve as the referring expressions.
-Furthermore, a dynamic feedback loop triggers color categorization (Step 2) when novel colors are detected, ensuring color attribute consistency.
-Step 4: quality evaluation. An evaluator agent assesses the annotation semantic accuracy and referential uniqueness of object-expression pairs, categorizing annotations as:
-  • “Yes.” The annotation is accurate and unique to the described objects.
-  • “No.” The annotation is inaccurate or non-unique, with a detailed explanation of the discrepancy.
-Annotations marked “Yes” proceed to Step 5.
-Annotations marked “No” activate specific feedback: expressions with semantic issues are returned to Step 3 (expression generation), while color-related inaccuracies are sent to Step 2 (color categorization) for refinement.
+Furthermore, a dynamic feedback loop triggers Step 2 (color categorization) when unseen colors are detected, ensuring color attribute consistency.
+Step 4: quality evaluation.
+A validation module assesses the semantic accuracy and uniqueness of each object-expression pair, classifying each annotation as either correct (accurately and uniquely identifying the described objects) or incorrect (inaccurate or ambiguous, with an explanation of the identified issue).
+Annotations deemed correct proceed to Step 5 (human verification), while incorrect annotations trigger targeted feedback: those with semantic errors are returned to Step 3 (expression generation), and color-related inaccuracies are redirected to Step 2 (color categorization).
+Figure 3: Object number distribution per expression in gRefCOCO [28] and RefDrone datasets.
+Figure 4: Object size distribution histograms in RefDrone and gRefCOCO [28] datasets.
 Step 5: human verification. Human annotators review annotation outputs in three tiers:
-  • Direct acceptance. Annotations satisfying all quality criteria are approved for the final dataset.
-  • Refinement required. Annotations with minor errors are corrected through human editing.
-  • Significant issues. Annotations with major errors or unclear descriptions are marked for complete reworking.
-Annotations with significant issues enter a feedback loop, returning to Step 4 (quality evaluation).
-If an annotation repeatedly fails to meet standards, it becomes a no-target sample.
-These feedback refinement loops ensure that all annotations meet our quality standards and that no-target expressions still relate to the image content.
-Each step utilizes LMMs through in-context learning with manually designed prompts and examples (see Supplementary Material).
-RDAgent requires only \$0.0539 pre expression in GPT4-o API cost and reduces human annotation effort by 88% compared to fully manual annotations.
-This cost-performance balance makes the framework scalable for large-scale datasets.
-Figure 4: Object number distribution per expression in gRefCOCO [22] and RefDrone datasets.
+  • Direct acceptance. Annotations satisfying all criteria are approved for the final dataset.
+  • Refinement required. Minor errors are corrected through human editing.
+  • Significant issues. Poorly grounded or inconsistent annotations trigger full regeneration.
+Annotations with significant issues re-enter the pipeline at Step 4 (quality evaluation).
+If an annotation repeatedly fails, it is classified as a no-target sample.
+These feedback loops ensure all annotations meet quality standards and that no-target expressions are contextually relevant to the image.
+Each step utilizes LMMs through in-context learning with carefully designed task-specific prompts and examples (see Appendix A.8).
+The final annotation quality demonstrates 42% directly accepted samples, 47% requiring minor refinement, and 11% demanding full re-annotation.
+RDAnnotator reduces human annotation effort by 85%, decreasing annotation time from 7 to 1 minute per expression, while incurring only \$0.0539 per expression in GPT-4o API costs.
+This cost-performance balance makes the framework scalable for large-scale dataset creation while maintaining annotation quality.
 Figure 5: Word frequency visualization in RefDrone dataset.
@@
 The RefDrone dataset contains 17,900 referring expressions across 8,536 images, comprising 63,679 object instances in 10 categories.
-We maintain the original train, validation, and test splits from VisDrone2019-DET [49].
-The dataset features an average expression length of 9.0 words, with each expression referring to an average of 3.8 objects.
-Figure 1 gives several examples that relate to the three key features and challenges in RefDrone:
-1) Multi-scale and small-scale target detection.
-Figure 3 presents an analysis of object scale distribution in RefDrone, with (a) showing the overall distribution and (b) offering a comparison with the gRefCOCO [22] dataset.
-In the RefDrone dataset, small-scale objects constitute 31% of all instances, while large-scale objects account for 14%.
-The detailed scale distribution in RefDrone demonstrates greater variance compared to gRefCOCO, highlighting the multi-scale challenges in RefDrone.
-2) Multi-target and no-target samples.
-The RefDrone dataset includes 11,362 multi-target and 847 no-target expressions, with the number of referred targets ranging from 0 to 242.
-Figure 4 illustrates the target number distribution, revealing higher complexity in multi-target scenarios compared to gRefCOCO, where expressions typically refer to no more than two objects.
-3) Complex environment with rich contextual expressions.
-The images present inherent complexity including diverse viewpoints, varying lighting conditions, and complex background environments.
+We retain the original train, validation, and test splits from VisDrone2019-DET [60].
+The average expression length is 9.0 words, and each expression refers to an average of 3.8 objects.
+Figure 1 highlights the three key challenges in RefDrone:
+1) Multi-target and no-target samples.
+Unlike conventional REC datasets [54, 36] that primarily focus on single-object references, RefDrone includes 11,362 multi-target and 847 no-target expressions.
+The number of referred targets per expression ranges from 0 to 242.
+Figure 4 illustrates the target number distribution, revealing higher complexity in multi-target scenarios compared to gRefCOCO, where the majority of expressions contain references to only one or two objects.
+2) Multi-scale and small-scale target detection.
+Figure 4 presents the challenges of scale distribution: small-scale objects ($<32^{2}$ pixels) make up 31% of instances, normal-scale ($32^{2}$-$96^{2}$ pixels) 55%, and large-scale ($>96^{2}$ pixels) 14%.
+Compared to gRefCOCO [28], RefDrone has a notably higher variance in object scales, emphasizing the dataset’s multi-scale challenges.
+3) Complex environment with rich contextual reasoning.
+The images present inherent complexity, including diverse viewpoints, varying lighting conditions, and complex background environments.
 The referring expressions go beyond simple object attributes (e.g., color, category) and spatial relationships (e.g., ‘left’, ‘near’), incorporating rich object-object interactions (e.g., ‘the white trucks carrying livestock’) and object-environment interactions (e.g., ‘the white cars line up at the intersection’).
-Figure 5 presents word cloud visualizations of (a) complete expressions and (b) background terms, highlighting the diverse vocabulary used in descriptions.
+Figure 5 visualizes this linguistic richness through word clouds of (a) complete expressions and (b) background terms, highlighting the diverse vocabulary used in descriptions.
 3.4 Dataset comparison
 Table 1 compares RefDrone with existing REC datasets.
-RefDrone stands out by its average number of referred objects and its use of LMM for expression generation.
-The expressions offer richer contextual details compared to template-based or human-annotated expressions.
-While RIS-CQ [13] employs LLM for expression generation, it lacks visual content during the generation process.
-This often results in expressions that, despite being linguistically complex, may be disconnected from visual content.
-Similarly, RSVG [47], although it focuses on small targets, falls short in terms of expression quality.
+RefDrone stands out for its higher average number of referred objects and use of LMM for expression generation.
+Its expressions offer richer contextual details than template-based or human-annotated expressions [59].
+While RIS-CQ [19] uses LLM for expression generation, it lacks visual content, often producing linguistically complex but visually disconnected expressions.
+Similarly, RSVG [56] focuses on small targets but lacks expression quality.
 In contrast, RefDrone comprehensively presents three challenges, establishing it as a challenging benchmark in REC tasks.
@@
 Previous benchmarks primarily focus on image-level metrics, which are sufficient for single-target or few-target samples.
-However, these metrics fail to handle expressions referring to multiple target objects—potentially scaling up to 100 instances per expression in our task.
-To address this limitation, we introduce instance-level metrics that provide a more granular evaluation of a model’s capability to identify individual target objects.
+However, these metrics fail to handle expressions referring to multiple target objects—potentially scaling up to hundreds of instances per expression in our task.
+To address this limitation, these new metrics provide a more granular evaluation of a model’s capability to identify individual target objects.
 Instance-level metrics: Accinst.
 and F1inst.
-measure the accuracy of individual bounding box predictions.
+provide a fine-grained evaluation of the ability to precisely localize individual target objects.
 We compute the intersection over union (IoU) between the true bounding boxes and the predicted bounding boxes.
@@
 We calculate:
-$\text{Acc}_{inst.}=\frac{TP+TN}{TP+TN+FP+FN}$ (1)
-and
+$\text{Acc}_{inst.}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}},$ (1)
 […]
@@
 and F1img.
-evaluate the overall accuracy of predictions per image.
-An image is considered a true positive (TP) if all predictions match the true bounding boxes, otherwise it is a false positive (FP).
-For no-target samples, a prediction without any bounding box is a true negative (TN), otherwise it is a false positive (FP).
-We calculate Accimg.
-and F1img.
-using the same formulas as Accinst.
-and F1inst., respectively.
-4 Proposed method
-We introduce Number GroundingDINO (NGDINO) to address multi-target and no-target challenges based on GDINO [28].
-Our insight is that utilizing number information of referred objects enhances model performance on these challenges.
-As illustrated in Figure 6, we introduce three components: (1) a number prediction head, (2) number-queries, and (3) a number cross-attention module.
+evaluate the overall accuracy at the image level.
+An expression is a true positive (TP) only if all predictions exactly match the true bounding boxes.
+Any partial or extra prediction is a false positive (FP).
+For no-target samples, no prediction is counted as a true negative (TN), otherwise a false positive (FP).
+Image-level metrics use the same formulations as instance-level metrics but are applied to whole expressions rather than individual objects.
+4 NGDINO
+We introduce Number GroundingDINO (NGDINO) to address multi-target and no-target challenges in REC.
+Our key insight is that explicit numerical reasoning about target counts enhances model performance on these challenges.
+NGDINO builds upon GDINO [34] while introducing three components: (1) a number prediction head, (2) learnable number-queries with number-guided query selection, and (3) a number cross-attention module.
+NGDINO adopts GDINO’s dual-encoder-single-decoder architecture.
+The image encoder extracts visual features from input images, while the text encoder processes referring expressions.
+A feature enhancer facilitates cross-modal fusion, and a language-guided query selection module initializes detection queries.
+The decoder iteratively refines detection queries through self-attention, image cross-attention, and text cross-attention mechanisms, ultimately producing bounding box predictions.
+We mainly improve the decoder part, as highlighted by the yellow box in Figure 6.
+Number prediction head.
+We introduce a classification module to predict the number of referred objects $N_{pred}$.
+This prediction head consists of an FFN layer applied to the detection queries $Q_{det}$, followed by mean pooling:
+$N_{prob}=\text{softmax}~{}(~{}\text{MeanPool~{}(~{}FFN}~{}(Q_{det}))),$ (3)
+$N_{pred}=\text{argmax}~{}(N_{prob}),$ (4)
+where $Q_{det}\in\mathbb{R}^{B\times L_{d}\times D}$ with batch size $B$, detection query length $L_{d}$, and feature dimension $D$.
+Rather than directly predicting arbitrary counts, we quantize the prediction space into five categories $\mathcal{C}=\{0,1,2,3,4+\}$, where $4+$ represents four or more targets.
+This quantization addresses the long-tailed distribution of target counts in real-world data (which follows Zipfian distribution [2]).
 Figure 6: Architecture of a single decoder layer in Number GroundingDINO.
 Key modifications from GDINO [34] (highlighted in yellow box) include: (1) number prediction head (FFN) to predict target number, (2) number-queries selected through the predicted number, (3) number cross-attention between selected number-queries and detection-queries.
-4.1 NGDINO
-Following GDINO, our model employs a dual-encoder-single-decoder architecture.
-The framework consists of an image backbone for visual feature extraction, a text backbone for text feature extraction, a feature enhancer for cross-modal fusion, a language-guided query selection module for detection-query initialization, and a cross-modality decoder for box refinement.
-Within decoder layers, detection-queries are fed into a self-attention layer, an image cross-attention layer to combine image features, a text cross-attention layer to combine text features, and an FFN layer.
-We mainly improve the decoder part, highlighted by the yellow box in Figure 6.
-The number prediction head, structured similarly to the detection head with an FFN layer, predicts the number of referred objects from detection-queries.
-These detection-queries are used to predict the objects, containing number information.
-The number-queries are learnable embeddings designed to capture various numerical patterns.
-These queries are initialized randomly with dimensions (bs, length_nquery, ndim), where bs is the batch size, length_nquery represents the number of query lengths, and ndim denotes the feature dimension.
-The predicted number guides the selection of number-queries through the query selection in Algorithm 1.
-The selected number-queries have a fixed length of length_snquery.
-We implement number cross-attention in parallel with self-attention to integrate the number information.
-In the number cross-attention, the selected number-queries serve as keys and values, while detection-queries serve as queries.
-The outputs are added to the self-attention features.
-4.2 Loss function
-[…]
-""" Input: detection_queries:(bs,length_dquery,ndim) number_queries:(bs,length_nquery,ndim) Output: selected_number_queries:(bs,length_snquery,ndim) """ number_feat = FFN(detection_queries) # (bs, length_dquery, 1) number_feat = number_feat.mean(dim=1) # (bs, 1) number = sigmoid(number_feat) # (bs, 1) index = slice(length_snquery*number, length_snquery*(number+1)) selected_number_queries = number_queries[index] # (bs, length_snquery, ndim)
-Algorithm 1 Query selection.
+Number-queries and number-guided query selection.
+We introduce a learnable set of number-queries $Q_{num}\in\mathbb{R}^{B\times L_{n}\times D}$, where $L_{n}$ is the length of number-queries.
+$Q_{num}$ is randomly initialized and trained to capture numerical patterns.
+Based on the predicted target number $N_{pred}$, we select a subset of number-queries $Q_{num}^{sel}$:
+$Q_{num}^{sel}=Q_{num}\big{[}:,\ L_{s}\cdot N_{pred}\ :\ L_{s}\cdot(N_{pred}+1),\ :\ \big{]},$ (5)
+where $L_{s}$ is the length of selected number-queries.
+This creates a direct mapping between the predicted number of objects and the specialized queries that encode patterns relevant to that cardinality.
+Number cross-attention module.
+We integrate the number information into the object detection process through a number cross-attention mechanism that operates in parallel with the self-attention in the decoder.
+In number cross-attention, detection queries $Q_{det}$ serve as queries (Q), while the selected number-queries $Q_{num}^{sel}$ serve as both keys (K) and values (V).
+The output of this module is added to the self-attention features before subsequent processing in the decoder layer.
+Training objective.
+We adopt GDINO’s bounding box supervision and add a Cross-Entropy loss for number prediction.
+Ablation studies determine the optimal parameters: selected number-query length $L_{s}=10$ and total number-query length $L_{n}=50$, matching our 5-category quantization scheme.
 5 Experiments
-5.1 Construction of RDAgent and NGDINO
-RDAgent.
-A straightforward method is to adapt RDAgent for REC tasks.
-This adaptation automates the process by replacing the object input with predictions from an object detector and removing the human verification step.
-For a fair comparison, we employ Faster R-CNN [34] as the object detector, which is trained on the VisDrone2019-DET [49] dataset.
-The agent component utilizes GPT-4o [^1] https://openai.com/index/hello-gpt-4o/.
-NGDINO.
-The NGDINO adopts a two-stage procedure to maintain training stability.
-First, we pre-train the number prediction head on the RefDrone dataset while initializing other components with GDINO [28] parameters.
-Then, we fine-tune the entire model.
-5.2 Implementation details
 We establish the benchmark with 13 representative models that can perform REC tasks, comprising 3 specialist models and 10 LMMs.
 The specialist models include MDETR [20], GLIP [25], and GDINO [34].
-For LMMs, we evaluate Shikra [5], ONE-PEACE [38], SPHINX-v2 [21], MiniGPT-v2 [3], Ferret [44], Kosmos-2 [31], Griffon [48], Qwen-VL [1], CogVLM [41], and LLaVA [24].
-Detailed model specifications are provided in the Supplementary Material.
-Zero-shot evaluation details.
-For zero-shot evaluation, we use the original model checkpoints as provided in their respective papers.
-The implementations and checkpoints for GLIP and GDINO are obtained from the MMDetection [4].
-Fine-tuning evaluation details.
-Our fine-tuning protocol maintains consistency across all experiments to ensure a fair comparison.
-We preserve the original learning strategies while excluding random crop augmentation due to its negative effect on position-sensitive samples.
-For LMMs, we employ the LoRA [11] fine-tuning strategy and follow the instruction tuning data structures.
-All fine-tuning experiments run for 5 epochs using 4 NVIDIA A100 GPUs.
-5.3 Experimental results
-Table 2: Experimental results of zero-shot baselines on RefDrone benchmark.
+For the LMMs, we evaluate Shikra [7], ONE-PEACE [47], SPHINX-v2 [27], MiniGPT-v2 [5], Ferret [53], Kosmos-2 [38], Griffon [57], Qwen-VL [3], CogVLM [50], and LLaVA [30].
+Detailed model specifications and implementation details are provided in the Appendix.
+Table 2: Experimental results of fine-tuning baselines on RefDrone benchmark.
 Top groups: specialist models.
@@
 Accimg.
-MDETR [14] 8.42 4.41 2.99 1.63 GLIP [19] 5.46 3.84 9.20 8.54 GDINO-T [28] 1.18 1.84 3.94 6.35 GDINO-B [28] 1.97 2.23 6.43 7.58 Shikra [5] 0.80 0.52 2.26 1.60 ONE-PEACE [38] 1.02 0.51 2.64 1.34 SPHINX-v2 [21] 1.59 0.80 4.61 2.36 MiniGPT-v2 [3] 2.69 1.36 6.38 3.29 Ferret [44] 3.18 1.62 8.48 4.43 Kosmos-2 [31] 8.06 4.20 8.64 4.52 Griffon [48] 9.16 4.81 16.77 9.18 Qwen-VL [1] 10.91 5.77 18.32 10.08 CogVLM [41] 15.38 8.33 30.73 18.15
-Zero-shot results.
-Table 2 presents the zero-shot evaluation results, assessing the models’ domain generalization capability.
-CogVLM [41] demonstrates state-of-the-art performance across multiple metrics.
-However, several advanced models, including Shikra [5], ONE-PEACE [38], and SPHINX-v2 [21], are limited to outputting only a single bounding box, due to constraints in their pre-training data or output strategy.
-This restriction significantly impacts their performance in multi-target scenarios.
-Table 3: Experimental results of fine-tuning baselines on RefDrone benchmark. Top groups: specialist models. Bottom groups: LMMs for REC tasks.
+MDETR [20] 32.60 19.49 19.17 10.81 GLIP [25] 24.23 14.86 16.92 13.29 GDINO-T [34] 30.50 19.02 29.65 21.07 NGDINO-T (Ours) 33.34 20.98 32.45 22.84 GDINO-B [34] 31.96 19.99 31.69 22.26 NGDINO-B (Ours) 34.44 21.76 34.01 23.89 MiniGPT-v2 [5] 4.97 2.74 13.56 8.97 LLaVA-v1.5 [30] 6.00 3.63 14.43 11.57 Qwen-VL [3] 14.14 7.61 20.10 11.17
+Table 3: Experimental results of zero-shot baselines on RefDrone benchmark. Top groups: specialist models. Bottom groups: LMMs for REC tasks.
 Methods F1inst.
@@
 Accimg.
-MDETR [14] 32.60 19.49 19.17 10.81 GLIP [19] 24.23 14.86 16.92 13.29 GDINO-T [28] 30.50 19.02 29.65 21.07 NGDINO-T (Ours) 33.34 20.98 32.45 22.84 GDINO-B [28] 31.96 19.99 31.69 22.26 NGDINO-B (Ours) 34.44 21.76 34.01 23.89 MiniGPT-v2 [3] 4.97 2.74 13.56 8.97 LLaVA [24] 6.00 3.63 14.43 11.57 Qwen-VL [1] 14.14 7.61 20.10 11.17 RDAgent (Ours) 58.14 41.13 37.07 23.54
+MDETR [20] 8.42 4.41 2.99 1.63 GLIP [25] 5.46 3.84 9.20 8.54 GDINO-T [34] 1.18 1.84 3.94 6.35 GDINO-B [34] 1.97 2.23 6.43 7.58 Shikra [7] 0.80 0.52 2.26 1.60 ONE-PEACE [47] 1.02 0.51 2.64 1.34 SPHINX-v2 [27] 1.59 0.80 4.61 2.36 MiniGPT-v2 [5] 2.69 1.36 6.38 3.29 Ferret [53] 3.18 1.62 8.48 4.43 Kosmos-2 [38] 8.06 4.20 8.64 4.52 Griffon [57] 9.16 4.81 16.77 9.18 Qwen-VL [3] 10.91 5.77 18.32 10.08 CogVLM [50] 15.38 8.33 30.73 18.15
+5.1 Experimental results
 Fine-tuning results.
-Table 3 illustrates the fine-tuning performance across different methods.
-The specialist model MDETR [14] shows strong performance in instance-level metrics, achieving comparable results to GDINO-B [28], but struggles in image-level understanding.
-Conversely, LMMs like Qwen-VL [1] demonstrate superior image-level comprehension (F1img.: 20.10%, Accimg.: 11.17%) while struggling with instance-level tasks (F1inst.: 14.14%, Accinst.: 7.61%).
-This performance disparity can be attributed to LMMs’ strong global image understanding capabilities but limited effectiveness in detecting small objects due to input resolution constraints.
-The proposed RDAgent achieves superior results compared to existing approaches, particularly in instance-level metrics, surpassing GDINO-B [28] by 26.18% in F1inst.
-and 21.14% in Accinst..
-The relatively smaller gains in image-level metrics (5.38% in F1img.
-and 1.28% in Accimg.) highlight the importance of instance-level metrics.
-However, RDAgent’s multi-step pipeline and reliance on ChatGPT responses result in longer processing time compared to end-to-end approaches.
-The proposed end-to-end method, NGDINO, consistently improves over the baseline GDINO [28] across both backbone architectures.
-Figure 7 shows the visualization results comparing NGDINO and baseline GDINO, demonstrating NGDINO’s effectiveness with multi-target samples.
+Table 3 illustrates fine-tuning performance across multiple methods on our RefDrone dataset.
+The consistently poor performance across all approaches highlights the inherent challenges of the proposed dataset.
+The specialist model MDETR [20] shows strong performance in instance-level metrics, achieving comparable results to GDINO-B [34], but struggles with image-level understanding.
+Conversely, LMMs like Qwen-VL [3] achieve superior image-level comprehension while struggling with instance-level tasks.
+This performance disparity can be attributed to LMMs’ effective global image understanding capabilities but limited effectiveness in detecting small objects due to input resolution constraints (detailed results by object scale are provided in Appendix Table 9).
+Our proposed NGDINO consistently outperforms the baseline GDINO [34] across both backbone architectures.
+Specifically, NGDINO-B achieves state-of-the-art performance, surpassing GDINO-B by significant margins in instance-level metrics (2.48% improvement in F1inst.
+and 1.77% in Accinst.) as well as image-level metrics (2.32% in F1img.
+and 1.63% in Accimg.).
+Zero-shot results.
+Table 3 presents the zero-shot evaluation results, assessing the models’ domain generalization capability.
+CogVLM [50] demonstrates state-of-the-art performance across metrics.
+However, several advanced models, including Shikra [7], ONE-PEACE [47], and SPHINX-v2 [27], are limited to outputting only a single bounding box, due to constraints in their pre-training data or output strategy.
+This restriction significantly impacts their performance in multi-target scenarios.
+Additional benchmark evaluations.
+As presented in Table 5, we evaluate our method on the gRefCOCO [28] benchmark with multi-target and no-target samples.
+The metric Pr@0.5 measures the percentage of predictions with $F1$ = 1 at IoU $\geq$ 0.5, while N-acc denotes the no-target accuracy.
+NGDINO-T demonstrates a notable improvement over GDINO-T, especially in the N-acc metric, achieving gains of 4.15% on test A and 1.39% on test B, which assess performance on no-target samples.
+In contrast, the performance gains in Pr@0.5 (0.36% and 0.74% on test A and B, respectively) are more modest.
+This can be partially attributed to the relatively simple structure of the multi-target queries in gRefCOCO, where expressions predominantly refer to only one or two objects.
+MDETR [20] achieves higher Pr@0.5 on the test A set.
+However, this is likely because MDETR tends to produce more outputs, resulting in lower N-acc performance due to an increased number of false positives on no-target examples.
+Table 5 demonstrates our method’s effectiveness on the aerial view RSVG [56] benchmark with multi-target samples, where NGDINO-T consistently outperforms previous SOTA methods.
 Table 4: Experimental results on gRefCOCO [28] dataset. Asterisk (*) denotes results reported in the gRefCOCO paper.
-Methods testA testB
-Pr@0.5 $\uparrow$ N-acc. $\uparrow$ Pr@0.5 $\uparrow$ N-acc. $\uparrow$
-MDETR* [14] 50.0 34.5 36.5 31.0
-UNINEXT* [43] 46.4 49.3 42.9 48.2
-GDINO-T [28] 45.69 79.02 44.83 76.69
-NGDINO-T 46.05 83.17 45.57 78.08
-To further validate the effectiveness of the proposed NGDINO, we conduct additional experiments on the gRefCOCO [22] dataset, which also includes multi-target and no-target samples (Table 4).
-NGDINO-T outperforms the baseline method GDINO-T, particularly in N-acc.
-metrics (by 4.15% in test A and 1.39% in test B), which evaluate accuracy on no-target samples.
-The limited improvements in Pr@0.5 metrics can be attributed to the relatively simple nature of multi-target samples in gRefCOCO, which typically contain only two objects per sample.
-Figure 7: Visualization results comparing NGDINO and GDINO [28], where red boxes indicate true positives and yellow boxes denote false negatives.
-Figure 8: Failure cases of NGDINO on RefDrone dataset. Red, green, and yellow boxes indicate true positives, false positives, and false negatives, respectively.
-Table 5: Ablation study on RefDrone dataset for the effectiveness of each component in NGDINO, with Swin-Tiny backbone. NPH denotes number prediction head. NCA denotes number cross-attention.
-NPH NCA Acc_$\_inst.$ Acc_$\_img.$ FPS 19.02 21.07 13.5 19.09 (0.07$\uparrow$) 21.18 (0.11$\uparrow$) 12.8 (0.7$\downarrow$) 19.51 (0.49$\uparrow$) 21.71 (0.53$\uparrow$) 12.9 (0.6$\downarrow$) 20.98 (1.96$\uparrow$) 22.84 (1.77$\uparrow$) 12.3 (1.2$\downarrow$)
-Table 6: Ablation study for the impact of varying selected number query length. Params indicates additional parameters introduced.
+Methods testA testB Pr@0.5 $\uparrow$ N-acc.
+$\uparrow$ Pr@0.5 $\uparrow$ N-acc.
+$\uparrow$ MDETR* [20] 50.0 34.5 36.5 31.0 UNINEXT* [52] 46.4 49.3 42.9 48.2 GDINO-T [34] 45.69 79.02 44.83 76.69 NGDINO-T 46.05 83.17 45.57 78.08
+Table 5: Experimental results on RSVG [56] dataset. Asterisk (*) denotes results reported in the EarthGPT [58] paper.
+Methods Pr@0.5 $\uparrow$ Pr@0.6 $\uparrow$ Pr@0.7 $\uparrow$ Pr@0.8 $\uparrow$ TransVG* [11] 72.41 67.38 60.05 49.10 MGVLF* [56] 76.78 72.68 66.74 56.42 EarthGPT* [58] 76.65 71.93 66.52 56.53 GDINO-T [34] 76.99 75.81 73.36 66.77 NGDINO-T 77.16 76.03 73.81 67.84
+Table 6: Ablation study of NGDINO components on the RefDrone dataset with Swin-Tiny backbone.
+Number Prediction Head Number Cross-Attention F1inst. Accinst. F1img. Accimg. FPS 30.50 19.02 29.65 21.07 13.5 30.62 19.09 29.90 21.18 12.8 31.25 19.51 30.73 21.71 12.9 33.34 20.98 32.45 22.84 12.3
+Table 7: Ablation study on the impact of varying selected number query length. Params indicates additional parameters introduced.
 Length F1inst. Accinst. F1img. Accimg. Params 1 31.63 19.80 31.23 22.12 1.58M 10 33.34 20.98 32.45 22.84 1.65M 100 32.06 20.10 32.23 22.81 2.34M
-Table 7: Ablation study for the effectiveness of RDAgent.
-Methods F1inst. Accinst. F1img. Accimg. ReCLIP [36] 24.62 14.04 11.58 6.15 GPT4-o 52.38 35.65 35.50 22.38 RDAgent 58.14 41.13 37.07 23.54
+Table 8: Experimental results of two-stage instance ranking methods on the RefDrone benchmark.
+Methods F1inst. Accinst. F1img. Accimg. ReCLIP [44] 24.62 14.04 11.58 6.15 GPT4-o 52.38 35.65 35.50 22.38 RDAnnotator 58.14 41.13 37.07 23.54
 5.2 Ablation studies
-Ablations of NGDINO components.
+Ablation study on NGDINO components.
 Table 6 presents the analysis of each component in NGDINO.
 The number prediction head alone shows a slight increase in Accinst.
-from 19.02% to 19.09%, indicating minimal impact of this auxiliary head on the referring tasks.
+from 19.02% to 19.09%, indicating the minimal impact of this auxiliary head on REC tasks.
 When introducing the number cross-attention without number selection, we observe a more substantial improvement, with Accinst.
-increasing from 19.02% to 19.51%.
-This improvement is attributed to the additional parameters introduced in the decoder.
-The significant improvement is achieved through the combination of both the number prediction task and number cross-attention components.
+increasing to 19.51%.
+This improvement can be partially attributed to the additional parameters introduced in the decoder.
+The most significant improvement is achieved through the combination of both the number prediction task and the number cross-attention components.
 This contribution increases Accinst.
-from 19.02% to 20.98%.
+to 20.98%.
 While the additional computational cost results in a minor decrease in inference speed (FPS), this trade-off yields a 1.96% increase in Accinst..
-Ablations of query length.
+Ablation study on query length.
 Table 8 analyzes the impact of varying the selected number query length.
 Utilizing a minimal query length of 1 lacks the capacity to capture complex numerical information.
-Conversely, extending the query length to 100 increases parameter count and introduces optimization difficulties, thereby adversely affecting performance.
+Conversely, extending the query length to 100 increases parameter count and computational overhead, potentially leading to optimization challenges.
 Through these experiments, we determine that a query length of 10 provides an optimal trade-off.
-Number prediction head performance.
+Performance on number prediction.
 The number prediction head achieves an overall number prediction accuracy of 53.5%.
 During the bounding box prediction stage, this accuracy is observed to be 22.84%.
-This performance gap demonstrates the effectiveness of number prediction head.
+This performance gap demonstrates the effectiveness of the number prediction head.
 Besides, the number prediction achieves a mean absolute error (MAE) of 0.51, suggesting high precision as predictions closely align with ground truth values.
-Ablations of RDAgent.
-Table 7 analyzes the effectiveness of the proposed RDAgent.
-The GPT4-o baseline is established using only Step 3 of RDAgent.
-We also compare against ReCLIP [36], a two-stage instance ranking method using CLIP [33] for REC tasks.
-All experiments utilize Faster-RCNN [34] for object detection.
-Results show that RDAgent consistently outperforms GPT4-o and ReCLIP across all metrics.
-However, the performance is partially limited by the Faster-RCNN object detector, which achieves only 18.0 mAP on the VisDrone2019-DET [49] dataset.
+Performance of two-stage instance ranking method.
+Table 8 analyzes our annotation framework RDAnnotator (without human verification) against alternative two-stage instance ranking methods for REC tasks.
+We compare: (1) GPT-4o [^1] https://openai.com/index/hello-gpt-4o/ using only Step 3 of RDAnnotator, and (2) ReCLIP [44], which employs CLIP [40] for instance ranking.
+All experiments utilize Faster-RCNN [41] as the object detector (18.0 mAP on VisDrone2019-DET [60]).
+Results show RDAnnotator’s ability to produce high-quality annotations.
+Furthermore, superior instance-level metrics over end-to-end methods reveal the promising potential of two-stage ranking approaches in multi-target challenges.
+Figure 7: Some failure cases of NGDINO on RefDrone dataset. Red, green, and yellow boxes indicate true positives, false positives, and false negatives, respectively.
 5.3 Limitations
 While NGDINO addresses the multi-target and no-target challenges, several challenges remain.
-Figure 8 shows typical NGDINO failure cases, highlighting challenges in the RefDrone dataset.
-The examples show rich contextual expressions, challenging backgrounds, and small-scale object detection.
-These challenging cases reflect real-world applications and highlight the areas for future improvement.
+Figure 7 shows some typical NGDINO failure cases, highlighting challenges in RefDrone.
+The examples show rich contextual reasoning, challenging backgrounds, and small-scale object detection.
+These challenging cases reflect real applications and highlight areas for future improvement.
 6 Conclusion
 In this work, we introduce RefDrone, a challenging benchmark specifically designed for referring expression comprehension in drone scenes.
-The dataset is constructed using the proposed RDAgent, a semi-auto annotation framework leveraging a multi-agent system.
-Furthermore, we develop NGDINO to address the multi-target and no-target challenges in the RefDrone dataset.
+The dataset is constructed using the RDAnnotator framework, an efficient semi-automated annotation pipeline combining LMM-based modules and human oversight.
+RDAnnotator provides high-quality annotations while reducing human effort.
+Furthermore, we develop NGDINO to address the multi-target and no-target challenges in RefDrone.
+Experimental results demonstrate consistently low performance across existing approaches, underscoring the inherent difficulty of the RefDrone benchmark.
 In the future, we aim to further enhance NGDINO to address additional challenges presented by RefDrone.
-We also plan to expand the benchmark to include referring expression segmentation and referring expression tracking tasks.
-Supplementary Material
-We provide supplementary material for further analysis, organized as follows:
+We also plan to extend RefDrone to larger-scale scenarios and introduce additional tasks such as referring expression segmentation and referring expression tracking.
+We believe RefDrone will serve as a valuable benchmark for advancing research in drone-based REC tasks.
+Appendix A Appendix
+We provide the following appendices for further analysis:
   1. 1.
-    Color attribute analysis of RefDrone. (Appendix A)
+    Details of the baseline method. (Appendix A.1)
   2. 2.
-    Details of baseline method. (Appendix B)
+    Implementation details. (Appendix A.2)
   3. 3.
-    Additional dataset examples. (Appendix C)
+    Color attribute analysis of RefDrone. (Appendix A.3)
   4. 4.
-    Resutlts w.r.t different object scales. (Appendix D)
+    Additional dataset examples. (Appendix A.4)
   5. 5.
-    Prompts and examples in RDAgent. (Appendix E)
-Appendix A Dataset color attribute analysis
-In the RefDrone dataset, color attributes are present in 69% expressions.
-We initially use the HSV color space to define six primary colors: white, black, red, blue, green, and yellow.
-During the annotation process, RDAgent expands the color set by adding six more colors: orange, pink, grey, purple, brown, and silver.
-The distribution of these color terms across the expressions is illustrated in Figure 9.
-Figure 9: Distribution of color terms in RefDrone expressions.
-Appendix B Details of baseline methods
+    Results with respect to different object scales. (Appendix A.5)
+  6. 6.
+    Results on RefCOCO/+/g datasets. (Appendix A.6)
+  7. 7.
+    Visualization results analysis. (Appendix A.7)
+  8. 8.
+    Prompts and examples in RDAnnotator. (Appendix A.8)
+A.1 Details of baseline methods
 The details for each baseline method:
@@
     • GLIP: Swin-Tiny with BERT-Base,
-      pretrained on Object365.
+      pretrained on Objects365.
     • GDINO-T: Swin-Tiny with BERT-Base,
-      pretrained on Object365, GoldG, GRIT, V3Det.
+      pretrained on Objects365, GoldG, GRIT, V3Det.
     • GDINO-B: Swin-Base with BERT-Base,
-      pretrained on Object365, GoldG, V3Det.
+      pretrained on Objects365, GoldG, V3Det.
   • Large multimodal models:
@@
     • CogVLM: 7B parameters, grounding-specific version.
-    • SPHINX-v2, Griffon: 13B parameters.
-Appendix C Dataset examples
+    • SPHINX-v2: 13B parameters.
+    • Griffon: 13B parameters.
+A.2 Implementation details
+NGDINO implementation.
+The NGDINO adopts a two-stage procedure to maintain training stability.
+First, we pre-train the number prediction head on the RefDrone dataset while initializing other components with GDINO [34] parameters.
+We then fine-tune the entire model.
+We also evaluate it on the gRefCOCO [28] dataset, which includes no-target and limited multi-target samples.
+Additionally, we evaluate it on RSVG [56], an aerial-view dataset with multi-target challenges.
+Zero-shot evaluation details.
+For zero-shot evaluation, we use the original model checkpoints as provided in their respective papers.
+The implementations and checkpoints for GLIP and GDINO are obtained from MMDetection [6].
+Fine-tuning evaluation details.
+Our fine-tuning protocol maintains consistency across all experiments to ensure a fair comparison.
+We preserve the original learning strategies while excluding random crop augmentation due to its negative effect on position-sensitive samples.
+For LMMs, we employ the LoRA [17] fine-tuning strategy and follow the instruction tuning data structures.
+All fine-tuning experiments are run for 5 epochs using 4 NVIDIA A100 GPUs.
+A.3 Details of color categorization.
+We design a hybrid color extraction pipeline consisting of two main components: (1) a WideResNet-101 classifier [55] trained on HSV color space-based labels.
+The color labels are refined via manual validation, ensuring robustness to annotation noise and label ambiguity.
+(2) An LMM-based color consistency module that validates color predictions through structured prompts.
+The latter cross-checks the classifier’s outputs to mitigate ambiguities caused by illumination variance or partial occlusion.
+In the RefDrone dataset, color attributes are present in 69% of expressions.
+We initially use the HSV color space to define six primary colors: white, black, red, blue, green, and yellow.
+During the annotation process, RDAnnotator expands the color set by adding six more colors: orange, pink, grey, purple, brown, and silver.
+The distribution of these color terms across the expressions is illustrated in Figure 8.
+Figure 8: Distribution of color terms in RefDrone expressions.
+A.4 Dataset examples
 To provide a comprehensive understanding of our RefDrone dataset, we present representative examples in Figure 10.
 These samples demonstrate the three key challenges in our dataset, highlighting its real-world applicability.
-Appendix D The results of different object scales.
-We add small/medium/large instance accuracy metrics (ACCs, ACCm, ACCl) of RefDrone dataset in Table 8.
-Table 8: Fine-tuned results w.r.t. different object scales.
-Methods Accs Accm Accl Methods Accs Accm Accl GLIP 2.65 18.28 27.29 LLaVA 4.18 7.26 11.99 GDINO-T 8.56 20.66 33.99 Qwen-VL 0.73 10.60 30.18 NGDINO-T 10.84 23.38 40.13 RDAgent 32.54 47.97 47.17
-Appendix E Prompts and examples for RDAgent
-In this section, we provide the prompts and examples employed in RDAgent.
+A.5 The results of different object scales.
+We add small/medium/large instance accuracy metrics (ACCs, ACCm, ACCl) for the RefDrone dataset in Table 9.
+Table 9: Fine-tuned results with respect to different object scales.
+Methods Accs Accm Accl Methods Accs Accm Accl
+GLIP 2.65 18.28 27.29 LLaVA 4.18 7.26 11.99
+GDINO-T 8.56 20.66 33.99 Qwen-VL 0.73 10.60 30.18
+NGDINO-T 10.84 23.38 40.13 RDAnnotator 32.54 47.97 47.17
+A.6 Results on RefCOCO/+/g datasets.
+Since the RefCOCO, RefCOCO+, RefCOCOg datasets contain only one instance per expression, the proposed NGDINO leverages the number branch primarily to address multi-instance and no-instance scenarios.
+As a result, the performance of NGDINO is relatively similar to that of GDINO on these datasets.
+Table 10: Results on RefCOCO/+/g datasets.
+RefCOCO RefCOCO+ RefCOCOg
+TestA TestB TestA TestB Val Test
+MDETR 90.4 82.67 85.52 72.96 83.35 83.31
+GDINO-T 91.4 86.6 87.5 74.0 85.5 85.8
+NGDINO-T 91.5 86.5 87.8 74.7 85.3 85.8
+A.7 Visualization results compared NGDINO with baseline GDINO.
+Figure 9 shows visualization results comparing NGDINO and the baseline GDINO on the proposed RefDrone dataset, demonstrating NGDINO’s effectiveness with multi-target samples.
+Figure 9: Visualization results comparing (a) GDINO [34] and (b) NGDINO, where red boxes indicate true positives and yellow boxes denote false negatives.
+A.8 Prompts and examples for RDAnnotator
+In this section, we provide the prompts and examples employed in RDAnnotator.
 Table 11 presents the prompt construction process for expression generation (Step 3), which includes the system prompt and few-shot in-context learning examples.
@@
 Additionally, the system prompts for the feedback mechanism are presented in Table 14.
-Figure 10: Dataset samples from RefDrone.
-Table 9: Illustration of RDAgent’s prompt construction for expression generation (Step 3).
+Figure 10: Dataset examples from RefDrone.
+Table 11: Illustration of RDAnnotator’s prompt construction for expression generation (Step 3).
 Few-shot in-context-learning examples are from fewshot_samples.
-A representative example is shown in Figure 10
-messages = [ {"role":"system", "content": f"""As an AI visual assistant, your role involves analyzing a single image.
-You are supplied with three sentences that caption the image, along with additional data about specific attributes of objects within the image.
-This can include information about categories, colors, and precise coordinates.
-Such coordinates, represented as floating-point numbers that range from 0 to 1, are shared as center points, denoted as (x, y), identifying the center x and y.
-When coordinate x tends to 0, the object nears the left side of the image, shifting towards the right as coordinate x approaches 1.
-When coordinate y tends to 0, the object nears the top of the image, shifting towards the bottom as coordinate y approaches 1.
-Your task is to classify the provided objects based on various characteristics, while also substantiating your classification.
-This classification should be thoroughly justified, with criteria including but not limited to relationships or relative locations of objects.
-To refer to a specific object, use the provided coordinates directly.
-Base your classification justifications on direct observations from the image, avoiding any hypothesizing or assumptions."""} ] for  sample in fewshot_samples:
-       messages.append({"role":"user", "content":sample[‘context’]})
-messages.append({"role":"assistant", "content":sample[‘response’]} )
+A representative example is shown in Table 12
 […]
-Table 10: One example to illustrate the input and output in RDAgent Step 3 (expression generation).
+Table 12: One example to illustrate the input and output in RDAnnotator Step 3 (expression generation).
 The top block shows the input contexts such as image, object, caption and color attributed.
@@
 Coordinates: [0.431, 0.051], [0.431, 0.15], [0.431, 0.101], [0.434, 0.256], [0.439, 0.523], [0.443, 0.813], [0.445, 0.902], [0.446, 0.982], [0.235, 0.43], [0.232, 0.486], [0.223, 0.674], [0.215, 0.904]
-Table 11: System prompts employed at different stages of RDAgent, replacing the default system messages shown in Table 9
+Table 13: System prompts employed at different stages of RDAnnotator, replacing the default system messages shown in Table 11
 System prompt: Step 1 scene understanding
@@
 Your task is to assess whether the given colors of specific objects match their appearance in the image.
-Respond with ”Yes” when the colors are appropriate.
-In cases where the colors are deemed inappropriate, respond with a concise ”No.” System prompt: Step 3 expression generation
+Respond with "Yes" when the colors are appropriate.
+In cases where the colors are deemed inappropriate, respond with a concise "No." System prompt: Step 3 expression generation
 As an AI visual assistant, your role involves analyzing a single image.
@@
 Your task is to confirm whether the description exclusively relates to the described objects without including any others in the visual.
-Respond ”yes” if it matches, or ”no” with an explanation if it does not.
-Table 12: RDAgent system prompts for feedback mechanism. Differences from Table 11 are highlighted
+Respond "yes" if it matches, or "no" with an explanation if it does not.
+Table 14: RDAnnotator system prompts for feedback mechanism. Differences from Table 13 are highlighted
 System prompt: Step 2 color categorization with feedback mechanism
@@
 Your task is to assess whether the given colors of specific objects match their appearance in the image.
-Respond with ”Yes” when the colors are appropriate.
-In cases where the colors are deemed inappropriate, respond with a concise ”No.” System prompt: Step 3 expression generation with feedback mechanism
+Respond with "Yes" when the colors are appropriate.
+In cases where the colors are deemed inappropriate, respond with a concise "No." System prompt: Step 3 expression generation with feedback mechanism
 As an AI visual assistant, your role involves analyzing a single image.

### v3

@@
 To address these challenges, we introduce RefDrone, an REC benchmark for drone scenes.
-RefDrone reveals three key challenges: 1) multi-target and no-target scenarios; 2) multi-scale and small-scale target detection; 3) complex environment with rich contextual reasoning.
-To efficiently construct this dataset, we develop RDAnnotator, a semi-automated annotation framework where specialized modules and human annotator collaborate through feedback loops.
+RefDrone reveals three key challenges: 1) multi-target and no-target scenarios; 2) multi-scale and small-scale target detection; 3) complex environments with rich contextual reasoning.
+To efficiently construct this dataset, we develop RDAnnotator, a semi-automated annotation framework where specialized modules and human annotators collaborate through feedback loops.
 RDAnnotator ensures high-quality contextual expressions while reducing annotation costs.
@@
 Comprehensive experiments with state-of-the-art REC methods demonstrate that NGDINO achieves superior performance on RefDrone, as well as on the general-domain gRefCOCO and remote sensing RSVG benchmarks.
-The data and code are completely available at https://github.com/sunzc-sunny/refdrone.
 1 Introduction
 Figure 1: Examples of the various challenges in RefDrone dataset.
-Drones/UAVs have become increasingly popular in our daily lives, serving both personal and professional purposes, such as entertainment, package delivery, traffic surveillance, and emergency rescue [1, 42, 45].
-Their ability to move freely and to observe broadly make them important platforms for Embodied AI applications [12, 35, 23, 14, 31].
-A crucial capability in Embodied AI is Referring Expression Comprehension (REC) [43, 48, 4, 49], which serves as a critical bridge between natural language understanding and visual perception.
-REC requires drones to localize specific objects in images based on natural language expressions.
-However, existing REC datasets primarily focus on ground-level perspectives, such as RefCOCO [54, 36] dataset.
-The application of REC in drone-based scenarios presents unique challenges.
-In this work, we introduce RefDrone, a challenging REC benchmark for drone scenes.
-RefDrone comprises 17,900 referring expressions annotated over 8,536 images, with 63,679 object instances.
-As illustrated in Figure 1, RefDrone poses three primary challenges that distinguish it from existing datasets: (1) multi-target and no-target samples, where expressions may refer to any number of objects, ranging from 0 to 242; (2) multi-scale and small-scale target detection, with 31% small objects and 14% large objects; and (3) complex environment with rich contextual reasoning, encompassing diverse viewpoints, lighting conditions, and intricate backgrounds.
-The referring expressions richly capture spatial relations, object attributes, and inter-object interactions.
-As summarized in Table 1, RefDrone offers greater diversity and complexity than previous REC benchmarks.
-We further evaluate 13 representative REC models using both fine-tuning and zero-shot settings.
-Notably, all models exhibit significantly poorer performance on RefDrone compared to standard REC datasets (e.g., GroundingDINO [34] yields 21.07 $Acc_{img.}$ on RefDrone vs.
-91.4 $Acc_{img.}$ on RefCOCO), highlighting the inherent difficulty and unique challenges posed by our benchmark.
-To efficiently construct RefDrone, we develop RDAnnotator, a semi-automated annotation pipeline for referring expression annotation in drone scenes.
-RDAnnotator leverages multiple specialized LMM-based modules, which collaborate within a feedback loop to generate and validate annotations.
-This tool reduces human involvement to quality control and minor adjustments, thereby significantly lowering annotation costs, while maintaining high-quality standards for complex expressions.
-RDAnnotator achieves a cost of just \$0.0539 per expression (GPT-4o API usage) and reduces human annotation time to under one minute per expression.
-This cost-efficiency makes RDAnnotator a scalable solution for large-scale dataset construction and can be readily adapted to other REC tasks.
-Furthermore, we propose Number GroundingDINO (NGDINO) to address the challenges of multi-target and no-target referring expressions.
-Our key insight is that explicitly utilizing the number information of referred objects enhances the handling of these scenarios.
-NGDINO includes three key components: (1) a number prediction head that estimates the count of target objects, (2) learnable number-queries that capture numerical patterns corresponding to varying object quantities, and (3) a number cross-attention module that fuses number queries with detection queries to inform object localization.
-Extensive experiments on both our RefDrone benchmark, as well as the general-domain gRefCOCO [28] and remote sensing RSVG [56] datasets, demonstrate that NGDINO substantially improves performance, particularly in the challenging multi-target and no-target cases.
-In summary, our contributions are listed as follows:
-  • RefDrone Benchmark: We introduce RefDrone, a comprehensive benchmark for referring expression comprehension in drone scenes.
-RefDrone poses three key challenges: multi-target/no-target scenarios, multi-scale/small-object detection, and complex contextual reasoning.
-We provide thorough baseline evaluations with 13 representative REC models.
-  • RDAnnotator Framework: We propose RDAnnotator, a semi-automated annotation framework that significantly reduces human effort while ensuring high-quality referring expressions.
-This framework is scalable for large-scale dataset construction and generalizable beyond drone imagery.
-  • NGDINO Method: We develop NGDINO, a novel approach that explicitly models the number of referred objects to address multi-target and no-target cases.
-NGDINO achieves state-of-the-art performance on RefDrone and shows consistent improvements on gRefCOCO and RSVG benchmarks.
+Drones/UAVs have become increasingly prevalent in both personal and professional applications, including entertainment, package delivery, traffic surveillance, and emergency rescue [adao2017hyperspectral, san2018uav, varga2022seadronessee].
+Their mobility and broad observational capabilities make them promising platforms for Embodied AI applications [fan-etal-2023-aerial, liu2023aerialvln, lee2024citynav, gao2024embodied, liu2024coherent].
+A fundamental capability in Embodied AI is Referring Expression Comprehension (REC) [sima2023embodied, Wang_2024_CVPR, cai2024bridging, wang2024polaris], which bridges natural language understanding and visual perception by localizing specific objects in images based on textual descriptions.
+However, existing REC datasets predominantly adopt ground-level perspectives (e.g., RefCOCO [refcoco]), leaving drone-based scenarios largely underexplored despite their unique challenges.
 Table 1: Comparison of REC datasets relevant to RefDrone.
 Avg.
-objects: mean objects per expression.
+objects: objects per expression.
 Avg.
-length: mean words per expression.
+length: words per expression.
 No target: expressions without referred objects.
-Expression type: method for expression generation.
+Expression type: expression generation method.
 Small target: percentage of small-scale objects.
-RefCOCO/+/g [54, 36] gRefCOCO [28] D3 [51] RIS-CQ [19] RSVG [56] RefDrone (Ours) Image source COCO [26] COCO [26] COCO [26] VG [22]+COCO [26] DIOR [24] VisDrone [60] Avg.
+RefCOCO/+/g [refcoco, refcocog] gRefCOCO [grefcoco] D3 [xie2024described] RIS-CQ [riscq] RSVG [zhan2023rsvg] RefDrone (Ours) Image source COCO [coco] COCO COCO VG [vg]+COCO DIOR [dior] VisDrone [visdrone] Avg.
 objects 1.0 1.4 1.3 3.6 2.2 3.8 Avg.
-length 3.6/3.5/8.4 4.9 6.3 13.2 7.5 9.0 No target Expression type Manual Manual Manual LLM (GPT3.5-t) Templated LMM (GPT4-o) Small target 0/0/0% 0.1% 6.3% - 17.2% 31.1%
+length 3.6/3.5/8.4 4.9 6.3 13.2 7.5 9.0 No target Expression type Manual Manual Manual LLM Templated LVLM Small target 0/0/0% 0.1% 6.3% - 17.2% 31.1%
+[…]
+To enable efficient dataset construction, we develop RDAnnotator, a semi-automated annotation pipeline for referring expression annotation in drone scenes.
+RDAnnotator leverages multiple specialized LVLM-based modules, which collaborate within a feedback loop to generate and validate annotations.
+By reducing human involvement to quality control and minor refinements, RDAnnotator achieves a cost of merely \$0.0539 per expression (GPT-4o API) and reduces human annotation time to under one minute per expression.
+This cost-efficiency makes RDAnnotator a scalable solution for large-scale dataset construction and can be readily adapted to other REC tasks.
+Furthermore, we propose Number GroundingDINO (NGDINO) to address multi-target and no-target challenges.
+Our key insight is that explicitly modeling the number of referred objects significantly enhances handling of these scenarios.
+NGDINO includes three components: (1) a number prediction head estimating target object counts, (2) learnable number-queries capturing numerical patterns for varying object quantities, and (3) a number cross-attention module fusing number queries with detection queries for enhanced localization.
+Extensive experiments on our RefDrone benchmark, as well as the general-domain gRefCOCO [grefcoco] and remote sensing RSVG [zhan2023rsvg] datasets, demonstrate that NGDINO achieves substantial improvements, particularly in multi-target and no-target cases.
+In summary, our contributions are listed as follows:
+  • RefDrone Benchmark: We introduce RefDrone, a comprehensive REC benchmark for drone scenes featuring three key challenges: multi-target/no-target scenarios, multi-scale/small-object detection, and complex contextual reasoning.
+We provide thorough evaluations of 26 representative REC models.
+  • RDAnnotator Framework: We propose RDAnnotator, a cost-effective semi-automated annotation framework that significantly reduces human effort while ensuring high-quality annotations.
+This framework is scalable for large-scale dataset construction and generalizable beyond drone imagery.
+  • NGDINO Method: We develop NGDINO, a novel approach explicitly modeling object counts to address multi-target and no-target cases.
+NGDINO achieves state-of-the-art performance on RefDrone with consistent improvements on gRefCOCO and RSVG.
 2 Related Works
 2.1 Referring expression understanding datasets
-Referring expression understanding identifies regions in visual content from natural language expressions.
-The main subtasks are Referring Expression Comprehension (REC), which outputs detection bounding boxes, and Referring Expression Segmentation (RES), which outputs segmentation masks.
-Datasets in this domain show increasing visual and linguistic complexity.
-Early datasets such as ReferIt [21] and RefCOCO [54] are pioneering but limited to simple expressions and single-target scenarios.
-Later datasets [32, 9, 28, 46, 19, 51] introduce more challenges.
-gRefCOCO [28] supports multi-target expressions.
-OV-VG [46] enables open-vocabulary queries.
-D3 [51] and RIS-CQ [19] offer more complex expressions.
-Domain-specific datasets [10, 8, 56, 37, 39] address specialized applications: RSVG [56] for remote sensing, RAVAR [37] for video action recognition, RIO [39] for affordance detection.
-However, drone scenes remain unexplored in REC, motivating our RefDrone benchmark.
-Table 1 compares RefDrone with other related REC benchmarks.
-Recent dataset creation approaches leverage large language models (LLMs) to reduce annotation costs.
-Works such as LLaVA [30] and Ferret [53] demonstrate the effectiveness of LLMs in generating instruction tuning data.
-Similarly, RIS-CQ [19] and RIO [39] employ LLMs to generate referring expressions.
-However, these approaches treat LLMs as text generators and may lack proper visual grounding.
-In contrast, we employ a semi-automated annotation framework with iterative feedback mechanisms throughout the annotation pipeline to reduce annotation costs and improve quality.
+Referring expression understanding identifies visual regions corresponding to natural language expressions.
+Subtasks include Referring Expression Comprehension (REC), which predicts bounding boxes, and Referring Expression Segmentation (RES), which generates pixel-level masks.
+Early benchmarks such as ReferIt [kazemzadeh2014referitgame] and RefCOCO [refcoco] were pioneering but mostly limited to single-target scenarios.
+Subsequent works introduced greater complexity.
+For instance, gRefCOCO [grefcoco] expanded to multi-target expressions, while D3  [xie2024described] and RIS-CQ [riscq] provided more descriptive and challenging language.
+However, these datasets typically involve a small, fixed number of targets per expression.
+Domain-specific extensions have also emerged, including video action recognition (RAVAR [peng2024ravar]) and affordance detection (RIO [qu2024rio]).
+In the aerial domain, RSVG [zhan2023rsvg] focuses on remote sensing data but uses expressions with limited contextual richness.
+AerialVG [Liu_2025_ICCV] focuses on using spatial relations for a single target, leaving the challenge of multi-target and no-target tasks.
+To fill this gap, our RefDrone benchmark provides a more complex and realistic setting by introducing complex multi-target scenarios with contextually rich expressions.
+Table 1 provides a detailed comparison with related benchmarks.
+Figure 2: The overview of the RefDrone annotation process with RDAnnotator.
+Multiple specialized LVLM-based modules collaborate both with each other and human annotators through iterative feedback loops to generate high-quality annotations.
 2.2 Referring expression comprehension methods
-REC methods can be broadly categorized into large multimodal models (LMMs) and specialist models.
-Large Multimodal Models [30, 53, 5, 7, 3, 38, 47, 27, 57, 50] have recently been applied to REC tasks as part of evaluating their broader visual-language understanding capabilities.
+REC methods can be broadly categorized into large vision language models (LVLMs) and specialist models.
+LVLMs [you2023ferret, chen2023minigpt, chen2023shikra, bai2023qwen, peng2023kosmos, lin2023sphinx, zhan2025griffon, wang2023cogvlm, jiang2025detect] have recently been applied to REC tasks as part of evaluating their broader visual-language understanding capabilities.
 These models leverage extensive referring instruction tuning data to achieve competitive performance without task-specific architectural designs.
-Despite their broad capabilities, LMMs consistently struggle with small object detection, primarily due to their inherent input resolution constraints.
-This constraint forces image downsampling, resulting in the loss of fine-grained visual details.
+To manage computational demands, LVLMs typically process downsampled images [llava] or employ visual token reduction strategies [bai2023qwen, wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels], which limits their ability to preserve fine-grained details.
+This constraint makes them less effective for grounding small objects, a crucial requirement in applications like aerial image analysis.
 Specialist models include two-stage and one-stage methods.
-Two-stage methods [18, 29, 16, 44, 15] typically approach REC as a ranking task: first generating region proposals through a detector, then ranking them based on their alignment with the language input.
-However, they are often criticized for slow inference speeds.
-In contrast, one-stage methods [20, 25, 33, 13, 52, 34] directly predict target regions guided by language input.
-These approaches leverage transformer to enable cross-modal interactions between visual and textual features.
-Among these methods, GroundingDINO (GDINO) [34] has gained widespread attention for its impressive results in REC tasks.
+Two-stage methods [hu2017modeling, liu2019learning, hong2019learning, reclip, han2024zero] typically approach REC as a ranking task: first generating region proposals through an object detector, then ranking proposals based on language-vision alignment.
+Despite achieving strong accuracy, two-stage pipelines suffer from slow inference.
+In contrast, one-stage methods [mdetr, glip, liu2023dq, gan2020large, yan2023universal, gdino] directly predict target regions guided by language input.
+These approaches leverage transformers to enable cross-modal interactions between visual and textual features.
+Among these, GroundingDINO (GDINO) [gdino] has gained widespread attention for its impressive results in REC tasks.
 Our work extends GDINO by incorporating explicit number modeling to handle multi-target and no-target scenarios, which are critical challenges in drone-based REC tasks.
 3 RefDrone benchmark
-Figure 2: The overview of the RefDrone annotation process with RDAnnotator.
-Multiple specialized LMM-based modules collaborate both with each other and human annotators through iterative feedback loops to generate high-quality annotations.
 3.1 Data source
-The RefDrone benchmark is built upon VisDrone2019-DET [60], a large-scale high-quality drone-captured object detection dataset.
-Source images are collected across multiple scenarios, illumination conditions, and flying altitudes.
-To ensure meaningful visual content, we filter out images with fewer than three objects and objects with bounding box areas below 64 pixels.
+The RefDrone benchmark is built upon VisDrone2019-DET [visdrone], a high-quality drone-captured object detection dataset.
+Images are collected across scenarios, illumination conditions, and flying altitudes.
+To ensure sufficient complexity for referring expression annotation, we filtered the dataset, retaining only images with at least three objects and excluding objects with bounding box areas smaller than 64 pixels.
 VisDrone2019-DET provides object categories and bounding box coordinates, which we convert to normalized center points (range 0-1).
-This approach reduces the input context for LMMs while preserving spatial information.
+This approach reduces the token count for LVLMs while preserving spatial relationships.
 3.2 RDAnnotator for semi-automated annotation
-RDAnnotator (referring drone annotator) is a semi-automated annotation framework that minimizes human effort while ensuring high quality.
-It integrates specialized modules powered by LMMs with human validation.
-As shown in Figure 2, RDAnnotator employs five structured steps.
+To construct our benchmark, we introduce RDAnnotator, a semi-automated annotation framework that balances annotation quality and scalability by integrating LVLM modules with a human-in-the-loop validation process.
+As illustrated in Figure 2, the framework operates in five steps.
 Step 1: scene understanding.
-A GPT-4o-based scene parsing module generates three diverse textual descriptions per image.
-These captions establish spatial layouts, object relationships, and contextual foundations to guide subsequent referring expression generation.
+A scene-parsing module, powered by GPT-4o, generates three diverse captions for each image.
+These captions establish a foundational context by detailing spatial arrangements and object relationships for subsequent referring expression generation.
 Step 2: color categorization.
-Color attributes are crucial discriminative features in referring expressions.
-We implement a hybrid color extraction pipeline, which combines low-level CNN-based classification (WideResNet-101 classifier [55]) with high-level LMM semantic reasoning, ensuring robust color attribution in challenging aerial conditions.
+To identify color attributes, which are crucial discriminative features in REC, we employ a hybrid pipeline.
+This pipeline combines a CNN-based classifier (WideResNet-101 [wideresnet]) with high-level LVLM-based semantic reasoning to ensure accurate color attribution despite challenging lighting and atmospheric conditions.
 Step 3: expression generation.
-To address the complexity of referring expression generation, we reformulate the referring expression generation task as an object grouping problem.
-The target of the module is to group semantically related objects and provide appropriate reasons for each group.
-The reasons serve as the referring expressions.
-Furthermore, a dynamic feedback loop triggers Step 2 (color categorization) when unseen colors are detected, ensuring color attribute consistency.
+We reformulate expression generation as an object grouping task.
+This module clusters semantically related objects and generates a textual justification for each grouping, which then serves as the referring expression.
+A dynamic feedback loop connects this step to color categorization (Step 2), triggering re-analysis if novel color terms are generated, thereby ensuring attribute consistency across the dataset.
 Step 4: quality evaluation.
-A validation module assesses the semantic accuracy and uniqueness of each object-expression pair, classifying each annotation as either correct (accurately and uniquely identifying the described objects) or incorrect (inaccurate or ambiguous, with an explanation of the identified issue).
-Annotations deemed correct proceed to Step 5 (human verification), while incorrect annotations trigger targeted feedback: those with semantic errors are returned to Step 3 (expression generation), and color-related inaccuracies are redirected to Step 2 (color categorization).
-Figure 3: Object number distribution per expression in gRefCOCO [28] and RefDrone datasets.
-Figure 4: Object size distribution histograms in RefDrone and gRefCOCO [28] datasets.
-Step 5: human verification. Human annotators review annotation outputs in three tiers:
+A validation module automatically assesses each generated object-expression pair for semantic accuracy and uniqueness.
+Correct annotations advance to human verification (Step 5).
+Incorrect annotations are routed back with targeted feedback: semantic errors are returned to expression generation (Step 3), while color inaccuracies are sent back to color categorization (Step 2).
+Figure 3: Object number distribution per expression in gRefCOCO [grefcoco] and RefDrone datasets.
+Figure 4: Object size distribution analysis.
+(a) Object size distribution in RefDrone dataset (small: $32^{2}=1024$ pixels, normal: 1024 to 9216 pixels, large: $96^{2}=9216$ pixels).
+(b) Object size histograms in RefDrone and gRefCOCO [grefcoco] datasets.
+Step 5: human verification. All generated annotations undergo a final human review, categorized into three tiers:
   • Direct acceptance. Annotations satisfying all criteria are approved for the final dataset.
-  • Refinement required. Minor errors are corrected through human editing.
-  • Significant issues. Poorly grounded or inconsistent annotations trigger full regeneration.
-Annotations with significant issues re-enter the pipeline at Step 4 (quality evaluation).
-If an annotation repeatedly fails, it is classified as a no-target sample.
-These feedback loops ensure all annotations meet quality standards and that no-target expressions are contextually relevant to the image.
-Each step utilizes LMMs through in-context learning with carefully designed task-specific prompts and examples (see Appendix A.8).
-The final annotation quality demonstrates 42% directly accepted samples, 47% requiring minor refinement, and 11% demanding full re-annotation.
-RDAnnotator reduces human annotation effort by 85%, decreasing annotation time from 7 to 1 minute per expression, while incurring only \$0.0539 per expression in GPT-4o API costs.
-This cost-performance balance makes the framework scalable for large-scale dataset creation while maintaining annotation quality.
+  • Refinement required. Annotations with minor errors are corrected through human editing.
+  • Significant issues. Annotations with poorly grounded or inconsistent content trigger full regeneration.
+This multi-level feedback system ensures high annotation fidelity.
+Expressions that consistently fail verification are designated as “no-target” samples, creating a set of contextually valid negative examples.
+Each step leverages LVLMs via in-context learning with carefully designed task-specific prompts and examples (see Appendix).
+The effectiveness of our framework is demonstrated by the annotation outcomes: 42% of samples were directly accepted, 47% required minor refinement, and only 11% were rejected for re-annotation.
+Ultimately, RDAnnotator reduces human annotation effort by 85% , decreasing the average time per expression from 7 minutes to 1 minute.
+This efficiency is achieved at a low API cost of \$0.0539 per expression, demonstrating the framework’s scalability for creating large-scale, high-fidelity REC datasets.
 Figure 5: Word frequency visualization in RefDrone dataset.
+Figure 6: Architecture of a single decoder layer in Number GroundingDINO.
+Key modifications from GDINO [gdino] (highlighted in yellow box) include: (1) a number prediction head (FFN) to estimate target count, (2) number-queries selected through the predicted number, (3) a number cross-attention between selected number-queries and detection-queries.
 3.3 Dataset analysis
-The RefDrone dataset contains 17,900 referring expressions across 8,536 images, comprising 63,679 object instances in 10 categories.
-We retain the original train, validation, and test splits from VisDrone2019-DET [60].
-The average expression length is 9.0 words, and each expression refers to an average of 3.8 objects.
-Figure 1 highlights the three key challenges in RefDrone:
+The RefDrone dataset comprises 17,900 referring expressions for 63,679 object instances across 8,536 images, spanning 10 categories.
+The dataset preserves the original train, validation, and test splits from VisDrone2019-DET [visdrone].
+On average, each expression is 9.0 words long and refers to 3.8 objects.
+RefDrone is characterized by three primary challenges, illustrated in Figure 1:
 1) Multi-target and no-target samples.
-Unlike conventional REC datasets [54, 36] that primarily focus on single-object references, RefDrone includes 11,362 multi-target and 847 no-target expressions.
-The number of referred targets per expression ranges from 0 to 242.
-Figure 4 illustrates the target number distribution, revealing higher complexity in multi-target scenarios compared to gRefCOCO, where the majority of expressions contain references to only one or two objects.
+In contrast to datasets like RefCOCO [refcoco] that focus on single-object references, RefDrone features a significant portion of complex queries, including 11,362 multi-target and 847 no-target expressions.
+The number of targets per expression ranges from 0 to 242.
+As shown in Figure 3, this distribution presents a greater challenge than that of gRefCOCO [grefcoco], where expressions typically refer to only one or two objects.
 2) Multi-scale and small-scale target detection.
-Figure 4 presents the challenges of scale distribution: small-scale objects ($<32^{2}$ pixels) make up 31% of instances, normal-scale ($32^{2}$-$96^{2}$ pixels) 55%, and large-scale ($>96^{2}$ pixels) 14%.
-Compared to gRefCOCO [28], RefDrone has a notably higher variance in object scales, emphasizing the dataset’s multi-scale challenges.
+The dataset exhibits a wide distribution of object scales (Figure 4): small objects ($32^{2}$ pixels) account for 31%, medium objects ($32^{2}$-$96^{2}$ pixels) for 55%, and large objects ($96^{2}$ pixels) for 14%.
+The high variance in object scales, particularly the prevalence of small objects, underscores multi-scale and small-scale target detection challenges.
 3) Complex environment with rich contextual reasoning.
-The images present inherent complexity, including diverse viewpoints, varying lighting conditions, and complex background environments.
-The referring expressions go beyond simple object attributes (e.g., color, category) and spatial relationships (e.g., ‘left’, ‘near’), incorporating rich object-object interactions (e.g., ‘the white trucks carrying livestock’) and object-environment interactions (e.g., ‘the white cars line up at the intersection’).
-Figure 5 visualizes this linguistic richness through word clouds of (a) complete expressions and (b) background terms, highlighting the diverse vocabulary used in descriptions.
-3.4 Dataset comparison
-Table 1 compares RefDrone with existing REC datasets.
-RefDrone stands out for its higher average number of referred objects and use of LMM for expression generation.
-Its expressions offer richer contextual details than template-based or human-annotated expressions [59].
-While RIS-CQ [19] uses LLM for expression generation, it lacks visual content, often producing linguistically complex but visually disconnected expressions.
-Similarly, RSVG [56] focuses on small targets but lacks expression quality.
-In contrast, RefDrone comprehensively presents three challenges, establishing it as a challenging benchmark in REC tasks.
+Images are captured in complex environments with diverse viewpoints, lighting, and dense backgrounds.
+Consequently, the referring expressions extend beyond simple attributes (e.g., color) and spatial relationships (e.g., ‘left of’) to describe complex object-object interactions (e.g., ‘the white trucks carrying livestock’) and object-environment interactions (e.g., ‘the white cars line up at the intersection’).
+This complexity is visualized in the word clouds in Figure 5.
+3.4 Comparison to existing datasets
+Table 1 situates RefDrone among existing REC datasets.
+Key distinguishing features of RefDrone are its high average number of targets per expression and the use of an LVLM annotation pipeline.
+This pipeline generates expressions with richer contextual details compared to those from template-based methods or human-only annotation [zhang2023multi3drefer].
+Although RIS-CQ [riscq] also uses an LLM for generation, its process is decoupled from visual content, leading to expressions that can be linguistically complex but visually ambiguous.
+While RSVG [zhan2023rsvg] targets small objects, the descriptive quality of its expressions is limited.
+In contrast, RefDrone comprehensively integrates these challenges, establishing it as a challenging benchmark in REC tasks.
 3.5 Evaluation metrics
-We introduce instance-level metrics extending traditional REC metrics to address multi-target challenges.
-Previous benchmarks primarily focus on image-level metrics, which are sufficient for single-target or few-target samples.
-However, these metrics fail to handle expressions referring to multiple target objects—potentially scaling up to hundreds of instances per expression in our task.
-To address this limitation, these new metrics provide a more granular evaluation of a model’s capability to identify individual target objects.
-Instance-level metrics: Accinst.
-and F1inst.
-provide a fine-grained evaluation of the ability to precisely localize individual target objects.
-We compute the intersection over union (IoU) between the true bounding boxes and the predicted bounding boxes.
-An IoU $\geq$ 0.5 is considered a true positive (TP), otherwise it is a false positive (FP).
-Unmatched true bounding boxes are counted as false negatives (FN).
-For no-target samples, a prediction without any bounding box is a true negative (TN), otherwise it is a false positive (FP).
-We calculate:
+To properly assess performance on multi-target expressions, we introduce instance-level metrics alongside traditional image-level ones.
+Standard image-level metrics, which treat the entire set of predictions for an expression as a single entity, are insufficient for our task.
+They cannot granularly penalize a model for missing individual objects within a large group.
+Our proposed instance-level metrics address this limitation by providing a fine-grained assessment of a model’s ability to localize each individual target.
+Instance-level metrics Accinst.
+and F1inst.: These metrics evaluate performance at the individual object level.
+We match each predicted bounding box to a ground-truth (GT) box.
+A prediction with an IoU $\geq$ 0.5 with a GT box is a true positive (TP).
+Unmatched predictions are false positives (FP), and unmatched GT boxes are false negatives (FN).
+For “no-target” samples, the absence of any prediction is a true negative (TN), while any prediction is an FP.
+Accuracy and F1-score are then computed as:
 $\text{Acc}_{inst.}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}},$ (1)
-[…]
-Image-level metrics: Accimg.
-and F1img.
-evaluate the overall accuracy at the image level.
-An expression is a true positive (TP) only if all predictions exactly match the true bounding boxes.
-Any partial or extra prediction is a false positive (FP).
-For no-target samples, no prediction is counted as a true negative (TN), otherwise a false positive (FP).
-Image-level metrics use the same formulations as instance-level metrics but are applied to whole expressions rather than individual objects.
+$\text{F1}_{inst.}=\frac{2\cdot\text{TP}}{2\cdot\text{TP}+\text{FP}+\text{FN}}.$ (2)
+Image-level metrics Accimg.
+and F1img.: These metrics assess performance on the entire expression, enforcing a stricter success criterion.
+For a given expression, a prediction is a true positive (TP) only if the set of predicted boxes perfectly matches the set of ground-truth boxes.
+Any mismatch (e.g., missing, extra, or inaccurate boxes) results in a false positive (FP).
+TN and FP for “no-target” samples are defined as above.
+The metrics are calculated using the same formulas but on an image-wide basis.
 4 NGDINO
-We introduce Number GroundingDINO (NGDINO) to address multi-target and no-target challenges in REC.
-Our key insight is that explicit numerical reasoning about target counts enhances model performance on these challenges.
-NGDINO builds upon GDINO [34] while introducing three components: (1) a number prediction head, (2) learnable number-queries with number-guided query selection, and (3) a number cross-attention module.
-NGDINO adopts GDINO’s dual-encoder-single-decoder architecture.
-The image encoder extracts visual features from input images, while the text encoder processes referring expressions.
-A feature enhancer facilitates cross-modal fusion, and a language-guided query selection module initializes detection queries.
-The decoder iteratively refines detection queries through self-attention, image cross-attention, and text cross-attention mechanisms, ultimately producing bounding box predictions.
-We mainly improve the decoder part, as highlighted by the yellow box in Figure 6.
+We introduce Number GroundingDINO (NGDINO), a novel architecture designed to address the challenges of multi-target and no-target referring expression comprehension (REC).
+Our core insight is that explicit numerical reasoning about target counts is crucial for accurately grounding such expressions.
+NGDINO builds upon the strong foundation of GDINO [gdino], inheriting its dual-encoder, single-decoder structure.
+Our contributions are concentrated within the decoder, as highlighted in Figure 6, where we introduce three key components: (1) a number prediction head, (2) learnable number-queries with number-guided query selection, and (3) a number cross-attention module.
 Number prediction head.
-We introduce a classification module to predict the number of referred objects $N_{pred}$.
-This prediction head consists of an FFN layer applied to the detection queries $Q_{det}$, followed by mean pooling:
-$N_{prob}=\text{softmax}~{}(~{}\text{MeanPool~{}(~{}FFN}~{}(Q_{det}))),$ (3)
-$N_{pred}=\text{argmax}~{}(N_{prob}),$ (4)
-where $Q_{det}\in\mathbb{R}^{B\times L_{d}\times D}$ with batch size $B$, detection query length $L_{d}$, and feature dimension $D$.
-Rather than directly predicting arbitrary counts, we quantize the prediction space into five categories $\mathcal{C}=\{0,1,2,3,4+\}$, where $4+$ represents four or more targets.
-This quantization addresses the long-tailed distribution of target counts in real-world data (which follows Zipfian distribution [2]).
-Figure 6: Architecture of a single decoder layer in Number GroundingDINO.
-Key modifications from GDINO [34] (highlighted in yellow box) include: (1) number prediction head (FFN) to predict target number, (2) number-queries selected through the predicted number, (3) number cross-attention between selected number-queries and detection-queries.
-Number-queries and number-guided query selection.
-We introduce a learnable set of number-queries $Q_{num}\in\mathbb{R}^{B\times L_{n}\times D}$, where $L_{n}$ is the length of number-queries.
-$Q_{num}$ is randomly initialized and trained to capture numerical patterns.
-Based on the predicted target number $N_{pred}$, we select a subset of number-queries $Q_{num}^{sel}$:
+To explicitly reason about the number of targets, we append a lightweight number prediction head to the decoder.
+This head takes the detection queries $Q_{det}$ as input, processes them through a Feed-Forward Network (FFN), and applies mean pooling to produce a probability distribution over target counts:
+$N_{prob}=\text{softmax}\left(\text{MeanPool}\left(\text{FFN}(Q_{det})\right)\right),$ (3)
+$N_{pred}=\text{argmax}(N_{prob}),$ (4)
+where $Q_{det}\in\mathbb{R}^{B\times L_{d}\times D}$ are the detection queries for a batch size $B$, query length $L_{d}$, and feature dimension $D$.
+To manage the long-tailed, Zipfian-like distribution of target counts [axtell2001zipf], we discretize the output space into five bins: $\mathcal{C}=\{0,1,2,3,4+\}$, where $4+$ represents four or more targets.
+This approach transforms the open-ended regression problem into a stable multi-class classification task.
+Number-guided query selection.
+To inject numerical priors into the decoding process, we introduce learnable number-queries, $Q_{num}\in\mathbb{R}^{B\times L_{n}\times D}$, where $L_{n}$ is the length of number-queries.
+These queries are trained to encode distinct numerical patterns.
+The predicted number, $N_{pred}$, then acts as an index to select a dedicated slice of these queries:
 $Q_{num}^{sel}=Q_{num}\big[:,\ L_{s}\cdot N_{pred}\ :\ L_{s}\cdot(N_{pred}+1),\ :\ \big],$ (5)
-where $L_{s}$ is the length of selected number-queries.
-This creates a direct mapping between the predicted number of objects and the specialized queries that encode patterns relevant to that cardinality.
+where $L_{s}$ is the length of number-queries per category. This mechanism creates a mapping between the predicted cardinality and queries encoding patterns.
 Number cross-attention module.
-We integrate the number information into the object detection process through a number cross-attention mechanism that operates in parallel with the self-attention in the decoder.
-In number cross-attention, detection queries $Q_{det}$ serve as queries (Q), while the selected number-queries $Q_{num}^{sel}$ serve as both keys (K) and values (V).
-The output of this module is added to the self-attention features before subsequent processing in the decoder layer.
+The selected number-queries, $Q_{num}^{sel}$, are fused with the detection queries, $Q_{det}$, via a number cross-attention module.
+This module operates in parallel with the self-attention layer within the decoder.
+The detection queries serve as the query (Q), while the selected number-queries serve as both key (K) and value (V).
+The output of this module is added to the output of the self-attention layer, enriching the detection queries with numerical context before they are passed to subsequent layers.
 Training objective.
-We adopt GDINO’s bounding box supervision and add a Cross-Entropy loss for number prediction.
-Ablation studies determine the optimal parameters: selected number-query length $L_{s}=10$ and total number-query length $L_{n}=50$, matching our 5-category quantization scheme.
+The model is trained end-to-end by augmenting the GDINO losses (for bounding box regression and label classification) with a Cross-Entropy loss for the number prediction task.
+Hyperparameters, including the selected number-query length ($L_{s}=10$) and total number-query length ($L_{n}=50$, representing 5 categories $\times$ 10 queries), were determined via ablation studies in the Appendix.
 5 Experiments
-We establish the benchmark with 13 representative models that can perform REC tasks, comprising 3 specialist models and 10 LMMs.
-The specialist models include MDETR [20], GLIP [25], and GDINO [34].
-For the LMMs, we evaluate Shikra [7], ONE-PEACE [47], SPHINX-v2 [27], MiniGPT-v2 [5], Ferret [53], Kosmos-2 [38], Griffon [57], Qwen-VL [3], CogVLM [50], and LLaVA [30].
-Detailed model specifications and implementation details are provided in the Appendix.
-Table 2: Experimental results of fine-tuning baselines on RefDrone benchmark.
-Top groups: specialist models.
-Bottom groups: LMMs for REC tasks.
-GDINO-T and GDINO-B denote GroundingDINO [34] with Swin-Tiny and Swin-Base backbones.
-The best results in each group are denoted with bold.
+We establish a comprehensive benchmark comprising 26 representative methods capable of performing REC tasks: 3 specialized REC methods, 7 task-specific LVLMs for REC, 12 general LVLMs with varying parameter scales, and 4 closed-source API models.
+To evaluate our proposed NGDINO, we conduct experiments on our RefDrone dataset and two public benchmarks: gRefCOCO [grefcoco] and RSVG [zhan2023rsvg].
+Table 2: Experimental results of zero-shot baselines on RefDrone benchmark. The best results in each group are denoted with bold.
+Categories Methods Params Time F1inst.
+Accinst.
+F1img.
+Accimg.
+Specialized REC Methods $\text{MDETR}_{\text{ResNet101}}$ [mdetr] 0.19B 2021/04/26 8.42 4.41 2.99 1.63 $\text{GLIP}_{\text{Swin-Tiny}}$ [glip] 0.15B 2021/12/06 5.46 3.84 9.20 8.54 $\text{GroundingDINO}_{\text{Swin-Tiny}}$ [gdino] 0.17B 2023/05/09 1.18 1.84 3.94 6.35 $\text{GroundingDINO}_{\text{Swin-Base}}$ [gdino] 0.23B 2023/05/09 1.97 2.23 6.43 7.58 Specialized LVLMs for REC Shikra [chen2023shikra] 7B 2023/07/03 0.80 0.52 2.26 1.60 $\text{ONE-PEACE}_{\text{Grounding}}$  [one_peace] 4B 2023/07/20 1.02 0.51 2.64 1.34 Kosmos-2 [peng2023kosmos] 1.6B 2023/10/30 8.06 4.20 8.64 4.52 $\text{CogVLM}_{\text{Grounding}}$ [wang2023cogvlm] 7B 2023/11/20 15.38 8.33 30.73 18.15 Griffon [zhan2025griffon] 13B 2023/12/06 9.16 4.81 16.77 9.18 Ferret [you2023ferret] 7B 2023/12/14 3.18 1.62 8.48 4.43 Rex-Omni [jiang2025detect] 3B 2025/10/15 54.06 37.10 43.90 28.52 General LVLMs 0 ~ 7B $\text{DeepSeek-VL2}_{\text{Tiny}}$ [wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels] 3B 2024/12/13 2.35 1.84 4.51 4.98 Qwen2.5-VL [bai2025qwen2] 3B 2025/02/20 40.00 25.06 38.22 23.89 Qwen3-VL [qwen3technicalreport] 4B 2025/10/15 48.68 32.31 47.78 32.28 General LVLMs 7B ~ 10B Qwen-VL [bai2023qwen] 7B 2023/08/12 10.91 5.77 18.32 10.08 MiniGPT-v2 [chen2023minigpt] 7B 2023/10/13 2.69 1.36 6.38 3.29 InternVL2.5 [internvl2_5] 8B 2024/12/05 0.58 1.14 1.79 4.2 Qwen2.5-VL [bai2025qwen2] 7B 2025/02/20 42.68 27.20 41.46 26.52 GLM 4.1V [vteam2025glm45vglm41vthinkingversatilemultimodal] 9B 2025/07/01 24.39 14.05 30.57 18.74 Ovis2.5 [lu2025ovis2] 9B 2025/08/19 4.24 2.20 14.56 8.39 $\text{MiMo-VL}_{\text{RL}}$ [coreteam2025mimovltechnicalreport] 7B 2025/08/21 19.75 11.22 30.24 18.79 InternVL3.5 [wang2025internvl3_5] 8B 2025/08/26 8.17 4.48 19.73 11.8 Qwen3-VL [qwen3technicalreport] 8B 2025/10/15 51.66 35.13 46.36 31.85 General LVLMs 10B ~ 30B SPHINX-v2 [lin2023sphinx] 13B 2023/11/17 1.59 0.80 4.61 2.36 $\text{DeepSeek-VL2}_{\text{Small}}$ [wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels] 16B 2024/12/13 34.95 21.34 36.15 22.81 DeepSeek-VL2 [wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels] 27B 2024/12/13 25.31 14.88 29.46 19.08 Qwen3-VL [qwen3technicalreport] 30B 2025/10/15 56.88 39.98 51.09 35.58 General LVLMs > 30B GLM 4.5V [vteam2025glm45vglm41vthinkingversatilemultimodal] 106B 2025/08/11 35.95 21.97 40.21 25.41 Qwen3-VL [qwen3technicalreport] 235B 2025/10/04 58.79 41.93 52.16 36.89 Closed-Source (APIs) Models DINO-XSeek [ren2024dino] 2025/03/11 54.47 37.46 46.11 30.14 Gemini 2.5 Pro [comanici2025gemini] 2025/03/25 3.45 1.91 8.11 4.90 Seed1.5-VL [guo2025seed1] 2025/05/12 43.52 27.84 37.27 23.03 Qwen3-VL-Plus [qwen3technicalreport] 2025/09/22 58.11 41.32 50.98 36.18 $\dagger$} State-of-the-art method.
+$\ddagger$} Models only predict a single bounding box, limiting multi-target performance.
+* GPT and Claude are excluded due to output bounding-box format failures.
+5.1 Zero-shot results.
+As presented in Table 2, we evaluate all models in a zero-shot setting to assess their generalization capabilities to the drone scenes.
+Overall performance.
+Among all models, the open-source Qwen3-VL (235B) [qwen3technicalreport] achieves state-of-the-art performance, with 58.79% F1inst., 41.93% Accinst., 52.16% F1img.
+, and 36.89% Accimg..
+These results surpass not only other open-source models but also the specialized closed-source API, DINO-XSeek [ren2024dino] (54.47% F1inst.), which is explicitly designed for REC tasks.
+Analysis of specialized methods.
+Specialized REC models (e.g., MDETR [mdetr], GLIP [glip], GroundingDINO [gdino]) show limited zero-shot transfer to RefDrone, with F1inst.
+generally below 10%.
+This likely stems from pre-training on general-domain datasets, which restricts their cross-domain adaptability.
+Task-specific LVLMs for REC show varied performance.
+Models limited to predicting a single bounding box, like Shikra [chen2023shikra], ONE-PEACE [one_peace], and SPHINX-v2 [lin2023sphinx], fail in multi-target scenarios, resulting in F1inst.
+scores below 2%.
+In contrast, recent Rex-Omni [jiang2025detect] achieves competitive performance (54.06% F1inst.
+/ 43.90% F1img.), approaching the state-of-the-art.
+Analysis of general LVLMs.
+A strong correlation is observed between model scale and zero-shot REC performance in general-purpose LVLMs.
+For instance, the Qwen3-VL series shows consistent improvement with scale: the 4B model achieves 48.68% F1inst., which increases to 51.66% for the 8B model and 56.88% for the 30B model.
+This scaling trend suggests that larger model capacity directly enhances localization abilities without task-specific tuning.
+A notable exception is DeepSeek-VL2 [wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels], where the small variant achieves better performance.
+Performance of closed-source APIs.
+Commercial models do not consistently outperform open-source alternatives.
+DINO-XSeek [ren2024dino] achieves 54.47% F1inst., lagging behind Qwen3-VL-235B.
+Other prominent APIs, including Gemini 2.5 Pro, yield low scores (F1inst.
+< 4%).
+Furthermore, models from the GPT and Claude families failed to produce outputs in the required bounding-box format.
+This disparity indicates that many leading commercial models are not optimized for fine-grained localization tasks like REC.
+Table 3: Results of fine-tuning baselines on RefDrone benchmark.
 Methods F1inst.
@@
 Accimg.
-MDETR [20] 32.60 19.49 19.17 10.81 GLIP [25] 24.23 14.86 16.92 13.29 GDINO-T [34] 30.50 19.02 29.65 21.07 NGDINO-T (Ours) 33.34 20.98 32.45 22.84 GDINO-B [34] 31.96 19.99 31.69 22.26 NGDINO-B (Ours) 34.44 21.76 34.01 23.89 MiniGPT-v2 [5] 4.97 2.74 13.56 8.97 LLaVA-v1.5 [30] 6.00 3.63 14.43 11.57 Qwen-VL [3] 14.14 7.61 20.10 11.17
-Table 3: Experimental results of zero-shot baselines on RefDrone benchmark. Top groups: specialist models. Bottom groups: LMMs for REC tasks.
-Methods F1inst.
-Accinst.
-F1img.
-Accimg.
-MDETR [20] 8.42 4.41 2.99 1.63 GLIP [25] 5.46 3.84 9.20 8.54 GDINO-T [34] 1.18 1.84 3.94 6.35 GDINO-B [34] 1.97 2.23 6.43 7.58 Shikra [7] 0.80 0.52 2.26 1.60 ONE-PEACE [47] 1.02 0.51 2.64 1.34 SPHINX-v2 [27] 1.59 0.80 4.61 2.36 MiniGPT-v2 [5] 2.69 1.36 6.38 3.29 Ferret [53] 3.18 1.62 8.48 4.43 Kosmos-2 [38] 8.06 4.20 8.64 4.52 Griffon [57] 9.16 4.81 16.77 9.18 Qwen-VL [3] 10.91 5.77 18.32 10.08 CogVLM [50] 15.38 8.33 30.73 18.15
-5.1 Experimental results
-Fine-tuning results.
-Table 3 illustrates fine-tuning performance across multiple methods on our RefDrone dataset.
-The consistently poor performance across all approaches highlights the inherent challenges of the proposed dataset.
-The specialist model MDETR [20] shows strong performance in instance-level metrics, achieving comparable results to GDINO-B [34], but struggles with image-level understanding.
-Conversely, LMMs like Qwen-VL [3] achieve superior image-level comprehension while struggling with instance-level tasks.
-This performance disparity can be attributed to LMMs’ effective global image understanding capabilities but limited effectiveness in detecting small objects due to input resolution constraints (detailed results by object scale are provided in Appendix Table 9).
-Our proposed NGDINO consistently outperforms the baseline GDINO [34] across both backbone architectures.
-Specifically, NGDINO-B achieves state-of-the-art performance, surpassing GDINO-B by significant margins in instance-level metrics (2.48% improvement in F1inst.
-and 1.77% in Accinst.) as well as image-level metrics (2.32% in F1img.
-and 1.63% in Accimg.).
-Zero-shot results.
-Table 3 presents the zero-shot evaluation results, assessing the models’ domain generalization capability.
-CogVLM [50] demonstrates state-of-the-art performance across metrics.
-However, several advanced models, including Shikra [7], ONE-PEACE [47], and SPHINX-v2 [27], are limited to outputting only a single bounding box, due to constraints in their pre-training data or output strategy.
-This restriction significantly impacts their performance in multi-target scenarios.
-Additional benchmark evaluations.
-As presented in Table 5, we evaluate our method on the gRefCOCO [28] benchmark with multi-target and no-target samples.
-The metric Pr@0.5 measures the percentage of predictions with $F1$ = 1 at IoU $\geq$ 0.5, while N-acc denotes the no-target accuracy.
-NGDINO-T demonstrates a notable improvement over GDINO-T, especially in the N-acc metric, achieving gains of 4.15% on test A and 1.39% on test B, which assess performance on no-target samples.
-In contrast, the performance gains in Pr@0.5 (0.36% and 0.74% on test A and B, respectively) are more modest.
-This can be partially attributed to the relatively simple structure of the multi-target queries in gRefCOCO, where expressions predominantly refer to only one or two objects.
-MDETR [20] achieves higher Pr@0.5 on the test A set.
-However, this is likely because MDETR tends to produce more outputs, resulting in lower N-acc performance due to an increased number of false positives on no-target examples.
-Table 5 demonstrates our method’s effectiveness on the aerial view RSVG [56] benchmark with multi-target samples, where NGDINO-T consistently outperforms previous SOTA methods.
-Table 4: Experimental results on gRefCOCO [28] dataset. Asterisk (*) denotes results reported in the gRefCOCO paper.
+GLIP-T [glip] 56.92 40.39 41.31 28.88 GDINO-T [gdino] 67.64 51.55 54.54 39.63 NGDINO-T (Ours) 71.11 55.52 56.51 41.20 GDINO-B [gdino] 69.75 53.95 56.95 41.81 NGDINO-B (Ours) 72.51 57.22 57.84 42.54
+5.2 Fine-tuning results.
+Performance on the RefDrone Dataset.
+Table 3 presents the fine-tuning performance across specialized REC methods for 50 epochs on RefDrone dataset.
+Our proposed NGDINO exhibits consistent gains over the GDINO baseline.
+Specifically, the Swin-Tiny models NGDINO-T surpasses GDINO-T by 3.47%, 3.97%, 1.97%, and 1.57% in F1inst., Accinst., F1img., and Accimg..
+Our NGDINO-B achieves the highest performance: 72.51% F1inst., 57.22% Accinst., 57.84% F1img.
+, and 42.54% Accimg., outperforming the state-of-the-art LVLM Qwen3-VL-235B by a significant margin.
+Beyond the performance advantage, NGDINO requires orders of magnitude fewer parameters than LVLMs, making it suitable for deployment on resource-constrained edge platforms such as drones.
+Table 4: Experimental results on gRefCOCO [grefcoco] dataset. Asterisk (*) denotes results reported in the gRefCOCO paper.
 Methods testA testB Pr@0.5 $\uparrow$ N-acc.
 $\uparrow$ Pr@0.5 $\uparrow$ N-acc.
-$\uparrow$ MDETR* [20] 50.0 34.5 36.5 31.0 UNINEXT* [52] 46.4 49.3 42.9 48.2 GDINO-T [34] 45.69 79.02 44.83 76.69 NGDINO-T 46.05 83.17 45.57 78.08
-Table 5: Experimental results on RSVG [56] dataset. Asterisk (*) denotes results reported in the EarthGPT [58] paper.
-Methods Pr@0.5 $\uparrow$ Pr@0.6 $\uparrow$ Pr@0.7 $\uparrow$ Pr@0.8 $\uparrow$ TransVG* [11] 72.41 67.38 60.05 49.10 MGVLF* [56] 76.78 72.68 66.74 56.42 EarthGPT* [58] 76.65 71.93 66.52 56.53 GDINO-T [34] 76.99 75.81 73.36 66.77 NGDINO-T 77.16 76.03 73.81 67.84
+$\uparrow$ MDETR* [mdetr] 50.0 34.5 36.5 31.0 UNINEXT* [yan2023universal] 46.4 49.3 42.9 48.2 GDINO-T [gdino] 45.69 79.02 44.83 76.69 NGDINO-T 46.05 83.17 45.57 78.08
+Performance on public benchmarks.
+To further validate the effectiveness of our improvements over GDINO in handling multi-target scenarios, we evaluate NGDINO on two benchmarks that contain multi-target samples: gRefCOCO [grefcoco] and RSVG [zhan2023rsvg].
+Models are fine-tuned for 5 epochs on both benchmarks.
+Table 4 presents results on gRefCOCO [grefcoco], which includes both multi-target and no-target samples.
+We adopt two metrics: Pr@0.5, measuring the percentage of predictions achieving $F1$ = 1 at IoU $\geq$ 0.5, and N-acc, denoting accuracy on no-target samples.
+NGDINO-T demonstrates improvements over GDINO-T in no-target detection, with N-acc gains of 4.15% and 1.39% on test A and test B, respectively.
+The improvements in Pr@0.5 are more modest (0.36% and 0.74% on test A and B), which can be attributed to the relatively simple multi-target structure in gRefCOCO, where expressions predominantly reference only one or two objects.
+While MDETR [mdetr] achieves higher Pr@0.5 on test A, this comes at the cost of lower N-acc due to its tendency to produce excessive outputs, leading to false positives on no-target samples.
+On the more complex aerial-view RSVG [zhan2023rsvg] benchmark (Table 5), NGDINO-T consistently outperforms prior state-of-the-art methods across all metrics.
+These results confirm that our proposed modifications effectively enhance performance on diverse multi-target REC tasks.
+Figure 7: Some failure cases of NGDINO on RefDrone dataset. Red, green, and yellow boxes indicate true positives, false positives, and false negatives, respectively.
+5.3 Ablation studies
+Analysis of NGDINO components.
+Table 6 presents the analysis of each component in NGDINO.
+With only the number prediction head, the model achieves improvements, improving F1inst.
+from 67.64% to 69.73% (+2.09%).
+This demonstrates that the auxiliary number prediction task enhances the model’s grounding capability.
+The number cross-attention mechanism alone yields more modest gains, improving F1inst.
+to 68.88% (+1.24%).
+This improvement can be partially attributed to the additional parameters introduced in the decoder.
+The most significant performance boost occurs when combining both components, where F1inst.
+reaches 71.11% (+3.47%).
+This substantial accuracy improvement comes at a negligible cost to efficiency, with only a minor decrease in inference speed (from 13.5 to 12.3 FPS).
+Number prediction accuracy.
+We evaluate the number prediction head’s performance.
+The head achieves a mean absolute error (MAE) of 0.21, indicating that its numerical predictions are highly precise and closely align with the ground truth counts.
+This strong performance in number prediction is a key factor in the model’s ability to handle multi-target expressions.
+The overall accuracy for predicting the correct number of instances is 75.3%.
+Table 5: Experimental results on RSVG [zhan2023rsvg] dataset. Asterisk (*) denotes results reported in the EarthGPT paper. Pr@0.5: percentage of predictions at IoU $\geq$ 0.5..
+Methods Pr@0.5 Pr@0.6 Pr@0.7 Pr@0.8 Pr@0.9 TransVG* [deng2021transvg] 72.41 67.38 60.05 49.10 27.84 MGVLF* [zhan2023rsvg] 76.78 72.68 66.74 56.42 35.07 EarthGPT* [zhang2024earthgpt] 76.65 71.93 66.52 56.53 37.63 GDINO-T [gdino] 76.99 75.81 73.36 66.77 49.96 NGDINO-T 77.16 76.03 73.81 67.84 51.39
 Table 6: Ablation study of NGDINO components on the RefDrone dataset with Swin-Tiny backbone.
-Number Prediction Head Number Cross-Attention F1inst. Accinst. F1img. Accimg. FPS 30.50 19.02 29.65 21.07 13.5 30.62 19.09 29.90 21.18 12.8 31.25 19.51 30.73 21.71 12.9 33.34 20.98 32.45 22.84 12.3
-Table 7: Ablation study on the impact of varying selected number query length. Params indicates additional parameters introduced.
-Length F1inst. Accinst. F1img. Accimg. Params 1 31.63 19.80 31.23 22.12 1.58M 10 33.34 20.98 32.45 22.84 1.65M 100 32.06 20.10 32.23 22.81 2.34M
-Table 8: Experimental results of two-stage instance ranking methods on the RefDrone benchmark.
-Methods F1inst. Accinst. F1img. Accimg. ReCLIP [44] 24.62 14.04 11.58 6.15 GPT4-o 52.38 35.65 35.50 22.38 RDAnnotator 58.14 41.13 37.07 23.54
-5.2 Ablation studies
-Ablation study on NGDINO components.
-Table 6 presents the analysis of each component in NGDINO.
-The number prediction head alone shows a slight increase in Accinst.
-from 19.02% to 19.09%, indicating the minimal impact of this auxiliary head on REC tasks.
-When introducing the number cross-attention without number selection, we observe a more substantial improvement, with Accinst.
-increasing to 19.51%.
-This improvement can be partially attributed to the additional parameters introduced in the decoder.
-The most significant improvement is achieved through the combination of both the number prediction task and the number cross-attention components.
-This contribution increases Accinst.
-to 20.98%.
-While the additional computational cost results in a minor decrease in inference speed (FPS), this trade-off yields a 1.96% increase in Accinst..
-Ablation study on query length.
-Table 8 analyzes the impact of varying the selected number query length.
-Utilizing a minimal query length of 1 lacks the capacity to capture complex numerical information.
-Conversely, extending the query length to 100 increases parameter count and computational overhead, potentially leading to optimization challenges.
-Through these experiments, we determine that a query length of 10 provides an optimal trade-off.
-Performance on number prediction.
-The number prediction head achieves an overall number prediction accuracy of 53.5%.
-During the bounding box prediction stage, this accuracy is observed to be 22.84%.
-This performance gap demonstrates the effectiveness of the number prediction head.
-Besides, the number prediction achieves a mean absolute error (MAE) of 0.51, suggesting high precision as predictions closely align with ground truth values.
-Performance of two-stage instance ranking method.
-Table 8 analyzes our annotation framework RDAnnotator (without human verification) against alternative two-stage instance ranking methods for REC tasks.
-We compare: (1) GPT-4o [^1] https://openai.com/index/hello-gpt-4o/ using only Step 3 of RDAnnotator, and (2) ReCLIP [44], which employs CLIP [40] for instance ranking.
-All experiments utilize Faster-RCNN [41] as the object detector (18.0 mAP on VisDrone2019-DET [60]).
-Results show RDAnnotator’s ability to produce high-quality annotations.
-Furthermore, superior instance-level metrics over end-to-end methods reveal the promising potential of two-stage ranking approaches in multi-target challenges.
-Figure 7: Some failure cases of NGDINO on RefDrone dataset. Red, green, and yellow boxes indicate true positives, false positives, and false negatives, respectively.
+Number Prediction Head Number Cross-Attention F1inst. Accinst. F1img. Accimg. FPS 67.64 51.55 54.54 39.63 13.5 69.73 53.90 56.29 41.08 12.8 68.88 52.89 54.98 39.74 12.9 71.11 55.52 56.51 41.20 12.3
 5.4 Limitations
-While NGDINO addresses the multi-target and no-target challenges, several challenges remain.
-Figure 7 shows some typical NGDINO failure cases, highlighting challenges in RefDrone.
-The examples show rich contextual reasoning, challenging backgrounds, and small-scale object detection.
-These challenging cases reflect real applications and highlight areas for future improvement.
+Despite its strong performance on multi-target and no-target scenarios, NGDINO has several limitations.
+As illustrated by the qualitative examples in Figure 7, failure cases typically arise from challenging scenarios inherent to the RefDrone dataset.
+These failures can be categorized into three primary sources: (1) complexity in the referring expression that demands sophisticated contextual reasoning; (2) cluttered backgrounds that camouflage target objects; and (3) inherent difficulties in detecting very small-scale objects.
+Addressing these challenging, real-world conditions remains an important direction for future work.
 6 Conclusion
-In this work, we introduce RefDrone, a challenging benchmark specifically designed for referring expression comprehension in drone scenes.
-The dataset is constructed using the RDAnnotator framework, an efficient semi-automated annotation pipeline combining LMM-based modules and human oversight.
-RDAnnotator provides high-quality annotations while reducing human effort.
+In this work, we introduce RefDrone, a challenging benchmark for referring expression comprehension in drone scenes.
+The dataset is built using RDAnnotator, an efficient semi-automated annotation pipeline that combines LVLMs with human-in-the-loop verification to ensure high-quality annotations.
+Our extensive experiments reveal a substantial performance drop for existing state-of-the-art methods when evaluated on RefDrone, underscoring the benchmark’s difficulty and exposing key limitations in current REC approaches.
 Furthermore, we develop NGDINO to address the multi-target and no-target challenges in RefDrone.
-Experimental results demonstrate consistently low performance across existing approaches, underscoring the inherent difficulty of the RefDrone benchmark.
 In the future, we aim to further enhance NGDINO to address additional challenges presented by RefDrone.
-We also plan to extend RefDrone to larger-scale scenarios and introduce additional tasks such as referring expression segmentation and referring expression tracking.
+We also plan to extend RefDrone to larger-scale scenarios and introduce additional tasks such as referring expression segmentation and tracking.
 We believe RefDrone will serve as a valuable benchmark for advancing research in drone-based REC tasks.
@@
 We provide the following appendices for further analysis:
-  1. 1.
-    Details of the baseline method. (Appendix A.1)
-  2. 2.
+  • Details of the baseline methods. (Appendix A.1)
   • Implementation details. (Appendix A.2)
-  3. 3.
-    Color attribute analysis of RefDrone. (Appendix A.3)
-  4. 4.
-    Additional dataset examples. (Appendix A.4)
-  5. 5.
-    Results with respect to different object scales. (Appendix A.5)
-  6. 6.
-    Results on RefCOCO/+/g datasets. (Appendix A.6)
-  7. 7.
-    Visualization results analysis. (Appendix A.7)
-  8. 8.
-    Prompts and examples in RDAnnotator. (Appendix A.8)
+  • Details of our color categorization pipeline.
+    (Appendix A.3)
+  • Results with respect to different object scales.
+    (Appendix A.4)
+  • Ablation study on query length. (Appendix A.5)
+  • Performance validation of the RDAnnotator framework. (Appendix A.6)
+  • Results on standard RefCOCO/+/g datasets.
+    (Appendix A.7)
+  • Additional examples from the RefDrone dataset.
+    (Appendix A.8)
+  • Prompts and examples used in RDAnnotator.
+    (Appendix A.9)
 A.1 Details of baseline methods
 The details for each baseline method:
-  • Specialist models:
   • MDETR: ResNet-101 with BERT-Base,
@@
     pretrained on Objects365, GoldG, V3Det.
-  • Large multimodal models:
-    • Kosmos-2: 1.6B parameters.
-    • ONE-PEACE: 4B parameters, visual grounding API.
-    • Shikra: 7B parameters, delta-v1 version.
-    • MiniGPT-v2: 7B parameters.
-    • LLaVA: 7B parameters, v1.5 version.
-    • Qwen-VL: 7B parameters.
-    • Ferret: 7B parameters.
-    • CogVLM: 7B parameters, grounding-specific version.
-    • SPHINX-v2: 13B parameters.
-    • Griffon: 13B parameters.
 A.2 Implementation details
 NGDINO implementation.
-The NGDINO adopts a two-stage procedure to maintain training stability.
-First, we pre-train the number prediction head on the RefDrone dataset while initializing other components with GDINO [34] parameters.
-We then fine-tune the entire model.
-We also evaluate it on the gRefCOCO [28] dataset, which includes no-target and limited multi-target samples.
-Additionally, we evaluate it on RSVG [56], an aerial-view dataset with multi-target challenges.
-Zero-shot evaluation details.
-For zero-shot evaluation, we use the original model checkpoints as provided in their respective papers.
-The implementations and checkpoints for GLIP and GDINO are obtained from MMDetection [6].
+We train NGDINO using a two-stage procedure to ensure stability.
+Stage 1: Pre-training.
+We initialize the model with weights from a pre-trained GDINO [gdino], freezing all components except for the number prediction head for 5 epochs.
+This new head is then pre-trained on the RefDrone dataset.
+Stage 2: End-to-end Fine-tuning.
+After the head is pre-trained, we unfreeze the entire model and fine-tune it on our target dataset.
+This staged approach prevents the randomly initialized prediction head from destabilizing the well-trained detector backbone during the initial phases of training.
+Zero-shot evaluation protocol.
+For all baseline models, we adhere to established evaluation practices to ensure fair comparisons.
+Specialized REC Methods: For GLIP [glip] and GDINO [gdino], we use the official model checkpoints and implementations provided within the MMDetection [mmdetection] framework.
+LVLMs: For a standardized and reproducible evaluation of LVLMs, we integrate our dataset into the VLMEvalKit framework [duan2024vlmevalkit].
+Within this framework, we benchmark each model using its official, recommended prompt structure to ensure optimal performance.
 Fine-tuning evaluation details.
-Our fine-tuning protocol maintains consistency across all experiments to ensure a fair comparison.
-We preserve the original learning strategies while excluding random crop augmentation due to its negative effect on position-sensitive samples.
-For LMMs, we employ the LoRA [17] fine-tuning strategy and follow the instruction tuning data structures.
-All fine-tuning experiments are run for 5 epochs using 4 NVIDIA A100 GPUs.
+All fine-tuning experiments are conducted within the MMDetection [mmdetection] framework on 8 NVIDIA A100 GPUs.
+To ensure a fair comparison, we apply a consistent protocol across all models.
+We follow the original learning strategies and hyperparameter settings for each model with one critical modification: we disable random crop data augmentation.
+This is because random cropping can remove crucial spatial context or the target objects themselves in our position-sensitive referring expressions, introducing label noise and degrading performance.
 A.3 Details of color categorization.
-We design a hybrid color extraction pipeline consisting of two main components: (1) a WideResNet-101 classifier [55] trained on HSV color space-based labels.
-The color labels are refined via manual validation, ensuring robustness to annotation noise and label ambiguity.
-(2) An LMM-based color consistency module that validates color predictions through structured prompts.
-The latter cross-checks the classifier’s outputs to mitigate ambiguities caused by illumination variance or partial occlusion.
-In the RefDrone dataset, color attributes are present in 69% of expressions.
-We initially use the HSV color space to define six primary colors: white, black, red, blue, green, and yellow.
-During the annotation process, RDAnnotator expands the color set by adding six more colors: orange, pink, grey, purple, brown, and silver.
-The distribution of these color terms across the expressions is illustrated in Figure 8.
+Color is a foundational attribute in the RefDrone dataset, present in 69% of all referring expressions.
+However, accurately identifying color is non-trivial due to challenges like illumination variance, occlusions, and semantic ambiguity (e.g., distinguishing "red" from "pink" or "orange").
+To address this, we designed a hybrid color extraction pipeline that combines the efficiency of a specialized classifier with the reasoning capabilities of an LVLM.
+The pipeline consists of two stages:
+(1) Classifier-based Proposal: A WideResNet-101 classifier [wideresnet] generates an initial color prediction.
+To create a high-quality training set for this classifier, we first generate labels programmatically using the HSV color space and then perform manual validation to correct noise and refine ambiguous cases.
+(2) LVLM Verification: An LVLM verifier then assesses the classifier’s output.
+Using structured prompts, it reasons about the visual evidence to confirm the prediction or correct it, effectively resolving ambiguities caused by lighting or partial visibility.
+The reliability of this hybrid approach enabled us to expand our vocabulary from an initial set of six primary colors (e.g., red, blue) to a more nuanced palette of twelve, including orange, pink, grey, and purple.
+The final distribution of these color terms is visualized in Figure 8.
 Figure 8: Distribution of color terms in RefDrone expressions.
-A.4 Dataset examples
+A.4 The results of different object scales.
+To provide a granular analysis of model robustness to scale variation, a critical challenge in the RefDrone benchmark, we evaluated a representative set of high-performing methods (ACCinst.
+> 10%) on objects categorized as small, medium, and large.
+The results, presented in Table 7, reveal a stark performance gap between LVLMs and Specialized REC methods, particularly on small objects.
+Notably, even LVLMs like Qwen3-VL-235B struggle, achieving only 26.54% Accs.
+In contrast, among the fine-tuned specialized REC methods, our NGDINO-B achieves 44.08% Accs, 62.59% Accm, and 68.20% Accl.
+Table 7: Performance comparison on small (Accs), medium (Accm), and large (Accl) objects.
+The models in the upper section are LVLMs evaluated in a zero-shot setting.
+The models in the lower section are specialized REC methods fine-tuned on the RefDrone training set.
+Methods Params Accs Accm Accl Accinst.
+Rex-Omni [jiang2025detect] 3B 25.55 43.03 50.31 37.10 Qwen2.5-VL [bai2025qwen2] 3B 13.52 28.59 32.44 25.06 Qwen2.5-VL [bai2025qwen2] 7B 12.19 33.36 35.80 27.20 Qwen3-VL [qwen3technicalreport] 4B 18.53 42.66 54.23 32.31 Qwen3-VL [qwen3technicalreport] 8B 20.05 43.64 50.32 35.13 Qwen3-VL [qwen3technicalreport] 30B 24.21 50.48 54.68 39.98 Qwen3-VL [qwen3technicalreport] 235B 26.54 51.60 55.79 41.93 $\text{MiMo-VL}_{\text{RL}}$ [coreteam2025mimovltechnicalreport] 7B 2.41 13.69 15.65 11.22 GLM 4.1V [vteam2025glm45vglm41vthinkingversatilemultimodal] 9B 1.25 16.94 29.84 14.05 GLM 4.5V [vteam2025glm45vglm41vthinkingversatilemultimodal] 106B 7.70 27.83 29.17 21.97 $\text{DeepSeek-VL2}_{\text{Small}}$ 16B 5.83 27.64 30.36 21.34 DeepSeek-VL2 [wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels] 27B 2.66 18.58 18.57 14.88 DINO-XSeek [ren2024dino] 24.54 46.52 62.51 37.46 Seed1.5-VL [guo2025seed1] 11.11 38.39 50.83 27.84 Qwen3-VL-Plus [qwen3technicalreport] 25.34 49.96 56.18 41.32 GLIP-T [glip] 0.15B 21.08 48.82 54.35 40.39 GDINO-T [gdino] 0.17B 38.56 56.69 66.13 51.55 GDINO-B [gdino] 0.23B 40.01 59.96 67.94 53.95 NGDINO-T (Ours) 0.18B 42.48 60.42 67.90 55.52 NGDINO-B (Ours) 0.24B 44.08 62.59 68.20 57.22
+A.5 Ablation study on query length.
+Table 8 analyzes the impact of varying the query length.
+A minimal query length of 1 lacks the capacity to capture complex numerical information.
+Conversely, extending the query length to 100 increases parameter count and computational overhead, potentially leading to optimization challenges.
+Through these experiments, we determine that a query length of 10 provides an optimal trade-off.
+Table 8: Ablation study on the impact of varying selected number query length. Params indicates additional parameters introduced.
+Length F1inst. Accinst. F1img. Accimg. Params 1 70.20 54.44 55.82 40.56 1.58M 10 71.11 55.52 56.51 41.20 1.65M 100 70.44 54.72 55.95 40.73 2.34M
+A.6 Performance of the RDAnnotator framework
+To validate the effectiveness of our proposed annotation framework RDAnnotator, we evaluate RDAnnotator when repurposed as a complete, two-stage method for REC.
+To ensure a fair comparison, all methods operate on an identical set of initial object proposals generated by a first-stage Faster-RCNN detector [fasterrcnn] (18.0 mAP).
+The core of the evaluation lies in the second stage, where each method uses the referring expression to rank these proposals and identify the target.
+We benchmark RDAnnotator against two strong alternative second-stage approaches: (1) GPT-4o, which represents a powerful, single-step LVLM reasoning approach, and (2) ReCLIP [reclip], a representative CLIP-based ranker that relies on embedding similarity.
+As shown in Table 9, RDAnnotator substantially outperforms both, validating the efficacy of its structured, multi-step reasoning process.
+Results underscore RDAnnotator’s suitability for generating high-fidelity REC annotations.
+Table 9: Experimental results of two-stage instance ranking methods on the RefDrone benchmark.
+Methods F1inst. Accinst. F1img. Accimg. ReCLIP [reclip] 24.62 14.04 11.58 6.15 GPT4-o 52.38 35.65 35.50 22.38 RDAnnotator 58.14 41.13 37.07 23.54
+A.7 Results on RefCOCO/+/g datasets.
+Since the RefCOCO, RefCOCO+, and RefCOCOg datasets contain only one instance per expression, the proposed NGDINO leverages the number branch primarily to address multi-instance and no-instance scenarios.
+As a result, the performance of NGDINO is relatively similar to that of GDINO on these datasets.
+Table 10: Results on RefCOCO/+/g datasets.
+RefCOCO RefCOCO+ RefCOCOg TestA TestB TestA TestB Val Test MDETR 90.4 82.67 85.52 72.96 83.35 83.31 GDINO-T 91.4 86.6 87.5 74.0 85.5 85.8 NGDINO-T 91.5 86.5 87.8 74.7 85.3 85.8
+A.8 Dataset examples
 To provide a comprehensive understanding of our RefDrone dataset, we present representative examples in Figure 9.
 These samples demonstrate the three key challenges in our dataset, highlighting its real-world applicability.
-A.5 The results of different object scales.
-We add small/medium/large instance accuracy metrics (ACCs, ACCm, ACCl) for the RefDrone dataset in Table 9.
-Table 9: Fine-tuned results with respect to different object scales.
-Methods Accs Accm Accl Methods Accs Accm Accl
-GLIP 2.65 18.28 27.29 LLaVA 4.18 7.26 11.99
-GDINO-T 8.56 20.66 33.99 Qwen-VL 0.73 10.60 30.18
-NGDINO-T 10.84 23.38 40.13 RDAnnotator 32.54 47.97 47.17
-A.6 Results on RefCOCO/+/g datasets.
-Since the RefCOCO, RefCOCO+, RefCOCOg datasets contain only one instance per expression, the proposed NGDINO leverages the number branch primarily to address multi-instance and no-instance scenarios.
-As a result, the performance of NGDINO is relatively similar to that of GDINO on these datasets.
-Table 10: Results on RefCOCO/+/g datasets.
-RefCOCO RefCOCO+ RefCOCOg
-TestA TestB TestA TestB Val Test
-MDETR 90.4 82.67 85.52 72.96 83.35 83.31
-GDINO-T 91.4 86.6 87.5 74.0 85.5 85.8
-NGDINO-T 91.5 86.5 87.8 74.7 85.3 85.8
-A.7 Visualization results compared NGDINO with baseline GDINO.
-Figure 9 shows visualization results comparing NGDINO and the baseline GDINO on the proposed RefDrone dataset, demonstrating NGDINO’s effectiveness with multi-target samples.
-Figure 9: Visualization results comparing (a) GDINO [34] and (b) NGDINO, where red boxes indicate true positives and yellow boxes denote false negatives.
-A.8 Prompts and examples for RDAnnotator
+Figure 9: Dataset examples from RefDrone.
+A.9 Prompts and examples for RDAnnotator
 In this section, we provide the prompts and examples employed in RDAnnotator.
@@
 Additionally, the system prompts for the feedback mechanism are presented in Table 14.
-Figure 10: Dataset examples from RefDrone.
 Table 11: Illustration of RDAnnotator’s prompt construction for expression generation (Step 3).